### v4

# Welfare Analysis in Dynamic Models

Abstract.
This paper introduces metrics for welfare analysis in dynamic models. We develop estimation and inference for these parameters even in the presence of a high-dimensional state space. Examples of welfare metrics include average welfare, average marginal welfare effects, and welfare decompositions into direct and indirect effects similar to Oaxaca (1973) and Blinder (1973).  We derive dual and doubly robust representations of welfare metrics that facilitate debiased inference. For average welfare, the value function does not have to be estimated. In general, debiasing can be applied to any estimator of the value function, including neural nets, random forests, Lasso, boosting, and other high-dimensional methods. In particular, we derive Lasso and Neural Network estimators of the value function and associated dynamic dual representation and establish associated mean square convergence rates for these functions.  Debiasing is automatic in the sense that it only requires knowledge of the welfare metric of interest, not the form of bias correction.  The proposed methods are applied to estimate a dynamic behavioral model of teacher absenteeism in Duflo et al. (2012) and associated average teacher welfare.
Keywords: Oaxaca-Blinder decomposition, value function, dynamic discrete choice, dynamic dual representation, average derivative, stationarity, finite dependence, double robustness
MIT, vchern@mit.edu
MIT, wnewey@mit.edu
University of California, Berkeley, vsemenova@berkeley.edu
1. Introduction
Dynamic considerations are important in applied work (Miller, 1984; Wolpin, 1984; Pakes, 1986; Rust, 1987; Hotz and Miller, 1993; Hotz et al., 1994; Keane and Wolpin, 1994; Aguirregabiria and Mira, 2002, 2007; Bajari et al., 2007). These considerations are captured by the value function, defined as the present discounted value of agents’ expected per-period utility. Prior work has focused on identifying utility parameters from the optimality of agents’ behavior and conducting inference on them. Building on this foundation, this paper targets a different yet complementary class of parameters that depend on the value function which we introduce below.
Welfare analysis is a central objective in economics. This paper develops identification, estimation, and inference results for welfare metrics in dynamic models. Examples of such metrics include the average value function and average marginal welfare effects, e.g., due to the change in observable time-invariant characteristics such as initial wealth. Most importantly, we decompose welfare effects into the direct effect through changes in utilities holding agents’ behavior fixed and the indirect effect from changes in agents’ behavior. This decomposition is analogous in spirit to Oaxaca-Blinder decomposition of outcome distributions yet different since welfare metrics are based on latent utilities rather than observed outcomes.
The economic motivation for the welfare decomposition we propose is drawn from various fields in applied economics, including labor economics, public finance, and health economics. In labor economics, Kitagawa (1955), Oaxaca (1973) and Blinder (1973) decompose the gender wage gap into structural and composition effects. In public finance, Chetty (2009) distinguishes between mechanical and behavioral  effects of taxation policies. In health economics, the work by Einav et al. (2010)  differentiates the direct effect of a health insurance price change from the indirect effect arising from changes in selection. For an example of cancer screening (Einav et al., 2020), the total welfare effect corresponds to the introduction of a new screening option, and the direct effect is its counterpart as if the frequency of screening were mandated. One of the paper’s contribution is to give identification, estimation, and inferential results for direct and indirect effects in dynamic models.
The next contribution of the paper is the dynamic dual representation of welfare metrics, which directly maps the per-period utility to the welfare metric of interest. For the case of average welfare, the dynamic dual representation reduces to a known expression of expected per-period utility. Furthermore, neither value function nor other dynamic object needs to be estimated for certain discrete choice models, such as those of   Rust (1987) and Hotz and Miller (1993).  For cases beyond average welfare, the dynamic dual representation involves backward discounting. In these cases,  we derive a doubly robust representation that facilitates consistent estimation if at least one of the two nuisance components, the value function or the dynamic dual representation,  is correct. Appendix C gives large sample properties of the proposed estimators based on Neyman-orthogonal moment equations, in which the dynamic dual representation plays a central role. This paper is the first result in the literature to leverage duality in dynamic models whose state distribution is strictly stationary.
Another contribution of the paper is to introduce a Lasso estimator for the value function, the dynamic dual representation, and associated mean square convergence rates in the setting of high-dimensional state variables. Specifically, the number of state variables may exceed the sample size, provided that only a small subset is relevant. We derive a novel least squares criterion that distinguishes our approach from previous IV-based methods. Adding an $\ell_{1}$ penalty to the sample analog of this criterion yields a Lasso estimator of the value function. In addition, we propose a neural network estimator and derive corresponding mean square convergence rates in the low-dimensional case. Although the results are presented for the value function, the methods apply to any fixed-point solution of an integral equation of the second kind – such as the Q-function in reinforcement learning – and do not require strict stationarity of the state distribution.
As an empirical application, we revisit the study of teacher absenteeism in the nonformal education centers (NFEs) of Rajasthan, India, originally analyzed by Duflo et al. (2012) (henceforth, DHR). First, we replicate their structural estimates using an infinite-horizon dynamic binary choice model as in Aguirregabiria and Mira (2002). Our estimates match those of DHR, which solves the finite-horizon model,  up to the third decimal place. This exercise demonstrates that our methodological results have broader empirical relevance: when the panel is sufficiently long, the distribution of the state vector can be treated as approximately stationary in the middle of the panel. Second, we incorporate  prior work history as a determinant of leisure utility. While these factors has strong predictive power, the structural estimates remain close to those reported by DHR. Finally, we estimate average welfare across different teacher groups determined by their score before they begin teaching.  Our welfare estimates exhibit interpretable patterns and are robust to the choice of regularized estimators.
1.1. Literature Review
A large body of work is dedicated to welfare analysis in discrete choice models with unobserved heterogeneity, as studied, e.g., in Bhattacharya (2015). As Bhattacharya (2024) discusses, welfare calculations are based on latent utilities rather than observed outcomes. For dynamic choice – the focus of this paper – important early references include Miller (1984), Wolpin (1984), and Pakes (1986). Within this group, a notable subclass of models includes those with the terminal action property, as in Hotz and Miller (1993) and Hotz et al. (1994), and, more broadly, models with finite dependence, as in Arcidiacono and Miller (2019). In Hotz and Miller (1993), structural parameters are identified via a regression problem that depends on conditional choice probabilities. The dynamic dual representation we derive extends this insight to the average value function and other welfare metrics.
Estimation of dynamic discrete choice models has received substantial attention; see, e.g., Aguirregabiria and Mira (2002, 2007); Bajari et al. (2007); Pesendorfer and Schmidt-Dengler (2008); Arcidiacono and Miller (2011); Arcidiacono et al. (2013a, b); Ackerberg et al. (2014); Aguirregabiria and Magesan (2018); Buchholz et al. (2021). Recent work by Adusumilli and Eckardt (2019) relaxes the terminal action property while allowing for an infinitely supported state space. Appendix F of the present paper outlines an automatic debiasing approach that targets nonlinear functionals of the value function and thus could be applicable to the model in Adusumilli and Eckardt (2019).
Last but not least, we contribute to the literature on estimating fixed points of integral equations of the second kind; see, e.g., Srisuma and Linton (2012). Prior work by Chen and Qi (2022) has shown that this is a special case of a well-posed NPIV problem and has developed value function estimators that achieve minimax-optimal rates in low-dimensional settings; see also Xia and Wainwright (2022) for related IV-based approaches. However, these methods may not readily extend to penalized estimators that remain consistent in high-dimensional settings. The least squares criterion we derive circumvents this limitation. We also give a least squares criterion for estimating the dynamic dual representation, that only depends on the welfare metric of interest, and so enables automatic debiasing like Chernozhukov et al. (2022) and Chernozhukov et al. (2024).
The paper is organized as follows. Section 2 presents the general framework and examples of welfare metrics. Section 3 decomposes the differences in average welfare in the spirit of Oaxaca and Blinder. Section 4 revisits Duflo et al. (2012) study of teacher absenteeism. Section 5 gives the dynamic dual and doubly robust representations for welfare metrics. Section 6 describes orthogonal estimating equations for average welfare and related averages. Least squares estimators of the value function and dynamic dual representation are presented in Section 7. Section 8 concludes. Appendix A considers the example of dynamic discrete choice. Appendix B gives proofs of the results in main text. Appendix C gives large sample properties of the estimators. Appendices DE give mean square rates for first-stage estimators. Appendix F extends the proposed method to nonlinear functionals of the value function. Appendix G demonstrates the method of Appendix F for dynamic binary choice.
2. Setup
We consider estimation and inference on welfare metrics that depend on the value function.  To define the value function,  let $X_{t},(t=0,1,...)$, denote a time series of observed state variables, which we assume to be a time-homogeneous, first-order Markov process with initial element $X:=X_{0}$.  The value function is determined by a per-period reward $\zeta_{0}(X)$, or, in other words,  expected utility in a single period conditional on the state $X$, which we assume to be identifiable. The value function $V_{0}(X)$ is the present discounted value of per-period rewards given the current state, satisfying
$V_{0}(X)=\sum_{t=0}^{\infty}\beta^{t}\mathrm{E}[\zeta_{0}(X_{t})\mid X],$ (2.1)
where $\beta\in[0,1)$ is a known discount factor. The value function satisfies the integral equation:
$V_{0}(X)=\zeta_{0}(X)+\beta\mathrm{E}[V_{0}(X_{+})\mid X],$ (2.2)
where $X$ is the current and $X_{+}$ is the next period element of the first-order Markov process (see Lemma B.1).  In what follows, we assume that the time series is strictly stationary. The welfare metrics we consider are linear functionals of the value function $V_{0}(X)$ having the form
$\delta_{0}=\mathrm{E}[w_{0}(X)V_{0}(X)],$ (2.3)
where $w_{0}(X)$ is some function of the state.
To give an example of per-period utility, we consider the dynamic discrete choice problem (Rust, 1987; Hotz and Miller, 1993; Aguirregabiria and Mira, 2002).  In each period, $(t=0,1,...),$ the agent chooses an action $j$ from a finite choice set $\mathcal{A}$.  The utility of choice $j$ in period $t$ is additively separable in a function of the current state $u(X_{t},j)$ and private shock $\varepsilon_{t}(j)$ and is given by
$\bar{u}(X_{t},j,\epsilon_{t})=u(X_{t},j)+\epsilon_{t}(j),\quad j\in\mathcal{A}.$
Here the sequence ${(X_{t},J_{t})}$ is strictly stationary, so the time index $t$ can be dropped.  The per-period reward is the expected utility $\zeta_{0}(x)$, obtained by taking the expectation over choices:
[…]
where ${\mathrm{P}}(J=j\mid X=x)$ is the probability that an agent chooses $j$ when $X=x$, and  $\mathrm{E}[\epsilon(j)\mid X=x,J=j]$ is the expectation of $\epsilon(j)$ under the choice $J=j$ and given $X=x$. For example, in a special case where the choice is binary and the private shocks are distributed as Gumbel,
$\zeta_{0}(x)=u(x,1)p_{0}(x)+u(x,0)(1-p_{0}(x))+H(p_{0}(x)),$ (2.4)
[…]
Within the discrete choice problems, our target parameter $\delta_{0}$ represents a welfare metric since $V_{0}(\cdot)$ is the expected value of an agent making optimal dynamic choices conditional on state $X$. Our first example is the expected value function where $w_{0}(X)=1$.
Example 2.1 (Average Welfare).
If $w_{0}(X)=1$, the parameter
$\delta_{0}=\mathrm{E}[V_{0}(X)]$ (2.5)
represents the unconditional expected value or welfare of making optimal dynamic choices.
We focus particularly on settings where the state variable include observable time-invariant heterogeneity in individual, per-period expected utility denoted by a vector $K$. Examples of $K$ include occupation type (Keane and Wolpin, 1997), gender and class grade of a child (Todd and Wolpin, 2006) and teacher test score (Duflo et al., 2012). Time invariant state variables could also represent observable heterogeneity in individual, per-period expected utility.     In this case, the state vector $X_{t}$ can be decomposed as $X_{t}=(S_{t},K),$  where $K_{t}=K$ does not vary over time.  Here first-order time homogeneity of $X_{t}$ implies that $(S_{t})_{t>0}$ is a first-order time-homogeneous Markov chain conditional on $K$.
Example 2.2 (Group Average Welfare).
When $K$ is discrete, taking on a finite number of values, the group average welfare is
$\delta_{0}=\mathrm{E}[1(K=k)V_{0}(X)]/{\mathrm{P}}(K=k)=\mathrm{E}[V_{0}(X)\mid K=k].$ (2.6)
In this case $w_{0}(X)=1(K=k)/{\mathrm{P}}(K=k)$.
Extending this example to differences in average welfare across groups is straightforward by differencing the parameter of interest in Example 2.2 across different values of $K$. In this example and the others, the weight $w_{0}(X)$ is unknown and will need to be estimated. The identification and estimation of $w_{0}(X)$ will be accounted for in the results that follow.
When $K$ represents an endowment of some resource it may be of interest to consider the welfare effect of changing the distribution of that endowment.
Example 2.3 (Average Policy Effect).
Let  $\pi(k)$ and $\pi^{*}(k)$ be the probability density (or mass) function of $K$ with respect to a base measure, corresponding to the actual data and a proposed policy shift.  The average policy effect from this shift is
$\delta_{0}$ $=\mathrm{E}[w_{0}(K)V_{0}(X)],\quad w_{0}(K)=[\pi^{*}(K)-\pi(K)]/\pi(K).$
This object differs from the policy effect of Stock (1989) in being the average effect of a policy on dynamic welfare rather than the average effect on some outcome variable.
For continuously distributed $K$ an effect of interest could be the average effect of changing $K$ on the value function.
Example 2.4 (Average Marginal Effect).
For continuously distributed $K$, the average derivative of the value function with respect to $K$,
$\delta_{0}=\mathrm{E}[\partial_{k}V_{0}(X)]=\mathrm{E}[\partial_{k}V_{0}(S,K)],$ (2.7)
measures the average change in welfare due to varying $K$.  Letting $f(K|S)$ denote the conditional PDF of $K$ given $S$, integration by parts gives equation (2.3) with  $w_{0}(X)=-\partial_{k}\ln f(K|S)$ as long as  $f(K|S)$ is equal to zero at the boundary of the support of $K$ conditional on $S$.  This parameter differs from the average derivative [^1] The average marginal welfare effects herein are different from the average marginal effects of Aguirregabiria and Carro (2024) in panel data setup, where marginal effects are taken with respect to unobserved unit heterogeneity. of Stoker (1986) in being a dynamic welfare effect rather than an outcome effect.
Example 2.5.
Let $\mathcal{X}$ be a set and $X$. Define
[…]
All of the above examples of target parameters fall in the following general theoretical  framework.   Let $Z$ denote a data vector that includes $X$ and $X_{+}$, and let $V$ denote a possible value function.  Also let $m(Z,V)$ denote a function of $Z$ and the function $V(\cdot)$ (i.e. $m(Z,V)$ is a functional of $V$.)  We consider parameters of the form
$\delta_{0}=\mathrm{E}[m(Z,V_{0})],$ (2.9)
where $\mathrm{E}[m(Z,V)]$ is linear in $V$. We will impose throughout that the expectation $\mathrm{E}[m(Z,V)]$ is mean square continuous as a function of $V$, meaning that there is a constant $C>0$ such that for all $V(X)$ with $\mathrm{E}[V(X)^{2}]<\infty$,
$|\mathrm{E}[m(Z,V)]|\leq C(\mathrm{E}[V(X)^{2}])^{1/2}.$ (2.10)
By the Riesz representation theorem mean square continuity of $\mathrm{E}[m(Z,V)]$ is equivalent to existence of a function $w_{0}(X)$ with $\mathrm{E}[w_{0}(X)^{2}]<\infty$ such that
$\mathrm{E}[m(Z,V)]=\mathrm{E}[w_{0}(X)V(X)]$ , UNKNOWN (2.11)
for all $V(\cdot)$ with $\mathrm{E}[V(X)^{2}]<\infty$. Here, we see that under mean square continuity, any parameter as in equation (2.9) can be represented as a linear function of the value function. There are many other potentially interesting examples of such welfare metrics. In the next section we consider decomposing differences in average welfare into direct and indirect components.
3. Decomposition of Differences in Average Welfare
To motivate our analysis, we present a simple running example in the context of a randomized controlled trial, in which a one-shot treatment assigned at time $t=0$ generates dynamic incentives. Suppose we aim to analyze welfare differences between the treated and control populations, denoted by $1$ and 00, respectively. In each group, the time-varying state variable is denoted by $S$, and the full state vector is $X=(S,K)$.
Let $V_{0}^{1}(s)=V_{0}(s,1)$ and $V_{0}^{0}(s)=V_{0}(s,0)$ denote the treated and control value functions. Define the treated average welfare as
$\delta_{\langle 1\mid 1\rangle}$ $=\mathrm{E}[V_{0}(X)\mid K=1]$ (3.1)
and the control average welfare as
$\delta_{\langle 0\mid 0\rangle}$ $=\mathrm{E}[V_{0}(X)\mid K=0].$ (3.2)
Both quantities are special cases of the group average welfare defined in Example 2.2. For $k\in\{1,0\}$, we have
[…]
where $\pi^{k}(s)$ denotes the probability distribution function of $S$ conditional on $K=k$. The counterfactual welfare metric
[…]
does not correspond to the group average welfare of any observable subpopulation. Instead, it is constructed by integrating the treated value function with respect to the stationary distribution of the control population. Provided that $\pi^{1}(s)$ and $\pi^{0}(s)$ share the same support, this parameter is well-defined.
The treatment-control welfare difference can be decomposed in the spirit of Kitagawa (1955), Oaxaca (1973), and Blinder (1973):
$\delta_{\langle 1\mid 1\rangle}-\delta_{\langle 0\mid 0\rangle}=\left[\delta_{\langle 1\mid 0\rangle}-\delta_{\langle 0\mid 0\rangle}\right]+\left[\delta_{\langle 1\mid 1\rangle}-\delta_{\langle 1\mid 0\rangle}\right].$ (3.5)
If treatment is randomly assigned, this welfare difference admits a causal interpretation.
The proposed decomposition has an intuitive interpretation when the treatment affects only per-period utilities and does not impact the state transition. We describe such empirical settings in Examples 3.1 and 3.2 below. In this case, the stationary distributions $\pi^{1}(s)$ and $\pi^{0}(s)$ differ solely due to agents making different optimal choices in the treated and control states, respectively. The first summand,
[…]
captures the difference in per-period utilities holding the distribution of states fixed at the control level. This term can be interpreted as the direct or mechanical effect. The second summand,
[…]
reflects changes in agents’ behavior that alter the distribution of states, holding the utilities. This term can be interpreted as the indirect or behavioral effect. The sum of the direct and behavioral effects gives the total welfare difference.
Example 3.1 (School attendance).
The study by Duflo et al. (2012) estimates a dynamic behavioral structural model in which teachers choose between working and taking leisure, with the goal of predicting their choices as a function of cash incentives for attendance. In contrast to their focus, we examine teacher factual and counterfactual welfare.
Let $K=1$ indicate assignment to a cash bonus for daily attendance. The state variable $S$ denotes the number of days attended since the beginning of the month, and the full state vector is $X=(S,K)$. The direct effect, $\delta_{\langle 1\mid 1\rangle}-\delta_{\langle 1\mid 0\rangle}$, captures the impact of the bonus on welfare through per-period utilities, holding attendance behavior fixed. The indirect effect, $\delta_{\langle 1\mid 0\rangle}-\delta_{\langle 0\mid 0\rangle}$, reflects changes in stationary attendance patterns (i.e., choice probabilities) induced by the treatment, holding per-period utilities fixed. A similar decomposition applies to the dynamic behavioral model studied in Todd and Wolpin (2006).
Example 3.2 (Breast cancer screening).
A standard approach to evaluating welfare in the context of cancer screening focuses on mortality outcomes. However, a broader perspective considers the costs of screening—such as time, resource use, and the psychological burden of false positives—which are naturally captured by welfare metrics based on latent utilities inferred from observed choices.
Let $K=1$ indicate assignment to a novel breast cancer screening technology. The state variable $S$ denotes the time elapsed since the most recent screening, and the full state vector is $X=(S,K)$. The direct effect, $\delta_{\langle 1\mid 1\rangle}-\delta_{\langle 1\mid 0\rangle}$, captures the impact of the new technology on welfare through per-period utilities, holding screening behavior fixed. The indirect effect, $\delta_{\langle 1\mid 0\rangle}-\delta_{\langle 0\mid 0\rangle}$, reflects changes in screening behavior induced by the treatment, holding utilities fixed.
Proposition 3.1 expresses the counterfactual welfare $\delta_{\langle 1\mid 0\rangle}$ as a linear functional of the treated value function that does not depend on the stationary distributions.
Proposition 3.1 (Decomposition of Differences in Average Welfare).
Suppose both density functions $\pi^{0}(s)$ and $\pi^{1}(s)$ have the same support of the state variable $S$. Then the counterfactual welfare $\delta_{\langle 1\mid 0\rangle}$ is a special case of equation (2.9) with
$m(Z,V)$ $=\dfrac{V(S,1)1\{K=0\}}{{\mathrm{P}}(K=0)}$ (3.8)
whose Riesz representation (2.11) holds with
$w_{0}(X)=w_{0}(S,K)=\dfrac{1\{K=1\}}{{\mathrm{P}}(K=0)}\dfrac{{\mathrm{P}}(K=0\mid S)}{{\mathrm{P}}(K=1\mid S)}.$ (3.9)
Recent work has extended classical decomposition methods to modern settings. Chernozhukov et al. (2022) derive the Riesz representer for the Average Treatment Effect on the Treated (ATET), and Vafa et al. (2024) provides an Oaxaca-Blinder decomposition of wage differences. Proposition 3.1 departs from these approaches by offering a decomposition of average welfare in dynamic models where welfare is based on latent utilities rather than observed outcomes.
We include the average counterfactual welfare $\delta_{\langle 1\mid 0\rangle}$ as Example 6.4 and discuss its estimation in Section 6. Notice that the parameters $\delta_{\langle 1\mid 1\rangle}$ and $\delta_{\langle 0\mid 0\rangle}$ are special cases of Example 2.2 with $k=1$ and $k=0$, respectively. Therefore, it is straightforward to extend this example to accommodate direct and indirect effects. We discuss the estimation of the counterfactual welfare measure and related effects further in Section 6.
Remark 3.1 (Overview of Related Literature).
(Kitagawa, 1955; Oaxaca, 1973; Blinder, 1973) pioneered the use of least squares methods for decomposing differences in average outcomes, such as wages. This approach was later extended to the distributional setting in DiNardo et al. (1996); Machado and Mata (2005); Fortin et al. (2011); Chernozhukov et al. (2013), with Chernozhukov et al. (2013) also providing inference methods. Other related contributions include Oaxaca and Ransom (1994); Kline (2011); Chernozhukov et al. (2013); Kline (2014); Guo and Basse (2021); Vafa et al. (2024). In the canonical wage example, groups $1$ and 00 correspond to men and women, respectively. The direct effect captures the difference in wage schedules faced by men and women and is often interpreted as a measure of discrimination or preferential treatment. The indirect effect reflects differences in job-related characteristics, such as skills or experience, and is typically interpreted as a composition or selection effect.
4. Duflo et al. (2012) revisited
We study how daily financial incentives affect teacher attendance in single-teacher nonformal education centers (NFEs) operated by the NGO Seva Mandir in tribal villages of Udaipur, Rajasthan, India. From 2003 to 2005, Duflo et al. (2012) conducted a randomized trial in which tamper-proof cameras recorded photos at school opening and closing. A school day was deemed valid if the two images were at least five hours apart and showed at least eight students present. At the end of each month, teachers earned a base salary of 500 Rupees (Rs) per month if they worked fewer than 10 days, plus a 50 Rs bonus for each additional day of work. The 10-day threshold for bonus activation created the dynamic incentive, which we focus on in this application [^2] We abstract from the firing threat, which appears negligible: no teacher was fired during the study period, even in cases of near-total absence. According to Duflo et al. (2012), Seva Mandir takes a long-term view in assessing teacher performance, which may explain the lack of dismissals..  Our dataset consists of daily attendance records for 57 teachers over an 18-month period, from January 1, 2004, to June 30, 2005, and a test score obtained from an exam before teaching. [^3] Following Duflo et al. (2012), the dataset includes only days when teachers actively choose between work and leisure. Holidays and weekends, though automatically counted toward teachers’ total days worked, are excluded from the estimation sample.
4.1. Replicating Duflo et al. (2012)
We revisit the dynamic behavioral model of Duflo et al. (2012), henceforth DHR. Let $t$ denote the day of the month, ranging from $1$ to $T=30$. On each day $t$, a teacher chooses between working ($j_{t}=1$) and taking leisure ($j_{t}=0$). On the final day of the month, the consumption utility is determined by the monthly paycheck
$\pi(d_{T})=500+50\cdot\max(d_{T}-10,0),$ (4.1)
where $500$ is the base salary, $d_{T}$ is the total number of days worked by day the final day $T$, and $50$ is the bonus.   For all days $t<T$, there is no consumption; utility accumulates only through leisure and is modeled as
$u(x_{t},0)$ $=\bar{x}_{t}^{\prime}\mu_{0}.$ (4.2)
Here, $\bar{x}_{t}\in\mathrm{R}^{p_{X}}$ is the state vector including a constant and, depending on the model, may include other observable teacher characteristics, and $\mu_{0}\in\mathrm{R}^{p_{X}}$ is a parameter to be estimated. The state vector is $X_{t}=(\bar{X}_{t},d_{t})$.   For example, in Model I of DHR, the leisure utility is assumed to be the same for all teachers, which corresponds to $u(x_{t},0)=\mu_{0}$ and $X_{t}=d_{t}$.  The per-period utilities are not discounted. The number of components in $\bar{x}_{t}$ is taken to be sufficiently small so that maximum likelihood estimators could be applied.
We consider a stylized dynamic binary choice model that is a special case of the one described in Section 2. Since teachers cannot be fired, the base salary should not affect their choice to work.  We decompose the total monthly bonus into daily payments.  In other words,  we consider the model where the utility of working ($j_{t}=1$) on day $t$ is given by
$u(d_{t},1)$ $=50\cdot\mu_{1}\cdot 1\{d_{t}-10\geq 0\},$ (4.3)
where the indicator function, referred to as “In the money” in DHR,  reflects the daily structure of the bonus in the stylized model.   The parameter $\mu_{1}$ converts consumption, measured in Rupees, into utility terms. The utility of leisure is the same as in equation (4.2). The stylized model (4.3)–(4.2) has the same monetary flow as the exact model (4.1)–(4.2),  up to discounting of the bonus. To abstract away from the finite horizon effects, we restrict the sample to calendar days 15, 16, and 17 of each month. In the absence of finite horizon restriction, the methods proposed in this paper accommodate a rich pattern of observable heterogeneity in (4.2) that would have been prohibitive for likelihood estimators.
Table 1  compares  exact DHR estimates (Columns (1)-(2))  to their infinite-horizon replicates (Columns (3)-(4)) for selected coefficients. First, the estimated bonus effect $0.039$ lies within the range 0.0160.0160.0160.0160.0490.0490.0490.049.  On average, the replicated standard errors exceed the exact ones by a factor ranging from $2.5$ to $3$.  Since only three days from each month are used, the drop in the precision is expected. Second, the sign of the teacher test score coefficient remains consistent and negative across specifications. This preliminary evidence suggests that the finite horizon effects are negligible, and the methods developed in this paper are applicable to this dataset.
4.2. Adding State Variables.
The utility of skipping work is difficult to quantify in Rupees. As mentioned by DHR, the choice to take leisure may depend on social norms and informal requests, past effort, or fatigue. The observables such as the teachers test score may not adequately capture teacher heterogeneity. Following DHR’s suggestion, [^4] On page 1225, footnote 16 Duflo et al. (2012) mention that teachers decision to work may be influenced by prior work history. we include a vector of prior work history in $\bar{x}_{t}$ as a state component. We anticipate that the work history is sufficiently rich to adequately represent teachers preference for leisure. In addition, we summarize this vector using a work streak variable defined as the number of consecutive days a teacher has worked without taking leisure. Formally, the work streak is
$a_{t+1}=(a_{t}+1)\cdot 1\{j_{t}=1\},\quad a_{0}=0.$ (4.4)
Table 2 shows the out-of-sample mean squared error (MSEs) for predicting a teacher’s decision to work, using models that sequentially expand the covariate set from Model I to Model V. Adding the teacher’s test score (Model II) yields minimal improvement. Including month dummies (Model III) results in a modest reduction in MSE. Incorporating the work streak (Model IV) leads to a substantial improvement: relative to Model III, the MSE falls by roughly 32% for Logit, 32% for Probit, and 35% for Random Forest. Finally, adding the full 90-day work history dummies (Model V) further improves prediction, halving the MSE relative to the baseline. These findings are consistent with DHR’s hypothesis that prior work history is relevant for capturing teacher’s decision to work, even though they do not reveal which aspect of the per-period utilities prior work history is relevant for.
Table 1. Results from the Exact (DHR) and Stylized Structural Models
DHR CNS
(1) (2) (3) (4)
“In the money” ($\mu_{1}$) 0.049 0.016 0.039 0.039
(0.001) (0.001) (0.003) (0.003)
Teacher Test Score $(\mu_{0,Score})$0.005 -0.011
(0.002) (0.005)
Notes: Table reports selected structural estimates for the dynamic behavioral model. Columns (1)-(2) replicate DHR. Columns (3)–(4) estimate a stylized model defined in equations (4.3)–(4.2). For the latter case, the sample consists of 57 teachers observed over 18 months as well observations from the middle days of each month (15, 16, and 17). The discount factor is $\beta=0.99$. See main text for details.
Table 2. Out-of-Sample Mean Squared Error for Predicting Decision to Work
Model Covariates (State Components) Logit Probit RF
(1) (2) (3)
I Days Worked 0.174 0.175 0.153
II $+$ Test Score 0.173 0.174 0.153
III $+$ Month Dummies 0.171 0.171 0.147
IV $+$ Work Streak 0.116 0.116 0.095
V $+$ Work History Dummies 0.062 0.063 0.069
Notes: This table reports the Mean Squared Error (MSE) for predicting teacher decision to work, using models that progressively add explanatory variables. Row 1 includes only the number of days worked in the current month. Row 2 adds the teacher’s test score. Row 3 introduces month dummies. Row 4 includes the work streak, defined as the number of consecutive days worked since the teacher’s last absence. Row 5 adds the full 90-day work history vector. Models are trained on 2004 data and evaluated on 2005 observations. Columns correspond to Logistic Regression (Logit), Probit Regression (Probit), and Random Forest (RF). See Table 4 for descriptive statistics of the variables.
4.3. Teacher Welfare.
We model the utility of leisure as a linear function
[…]
where $\bar{x}_{t}$ includes a constant, the teacher’s test score, and a vector of 90-day work history dummies. This specification allows for heterogeneity in leisure preferences based on recent work history. The structural parameters $(\mu_{1},\mu_{0})$ include 92 coefficients, which we estimate using Algorithm 1 in Appendix G. This algorithm requires conditional choice probabilities and several nuisance components, including parts of the value function and the dynamic dual representation. For choice probabilities, we consider both logistic regression and random forest estimators, which correspond to the columns of Table 3. The value function components are estimated via the Lasso-based method described in Section 7.
Table 3 presents our main empirical findings. Panel A reports the structural estimates, while Panels B through D present average welfare results. Columns (1)–(2) correspond to a simple leisure utility specification where $\bar{x}_{t}$ includes only a constant and test score (Model II in Table 2). Columns (3)–(4) correspond to the full specification that includes work history dummies (Model V). Structural estimates are similar across specifications and remain close to those in Duflo et al. (2012). As expected, standard errors increase in the richer specification due to the higher number of coefficients being estimated.
The primary objective of Duflo et al. (2012) is to assess the effect on student outcomes. Yet, it is equally important to evaluate the welfare of teachers whose professional lives were affecte by the experiment. According to Duflo et al. (2012), several teachers reported improved discipline and focus due to monitoring and cash incentive. However, their original study did not provide quantitative welfare measures. We fill this gap by reporting average welfare in Panels B, C, and D of Table 3.
To estimate welfare, we use the dual estimator described in Section 6. Instead of estimating value function, the estimator combines the structural parameters from Panel A and the estimated choice probabilities. Across all models, the average welfare ranges from 165 to 188 Rupees per period. These patterns are consistent with observed attendance rates (Table 4) and the negative sign of the test score coefficient in Panel A. In particular, lower-scored teachers receive consistently lower welfare estimates than their higher-scored counterparts. This suggests that policymakers may wish to consider setting a minimum test score threshold, making the bonus conditional on skill, or both.
Table 3. Welfare Estimates for Structural Model
Model A B
CCP Estimators Logit RF Logit RF
(1) (2) (3) (4)
Panel A: Estimates for Selected Coefficients in the Structural Model
“In the money” ($\mu_{1}$) 0.039 0.039 0.040 0.040
(0.003) (0.003) (0.005) (0.005)
Teacher Test Score $(\mu_{0,Score})$0.011 -0.011 -0.017 -0.017
(0.005) (0.005) (0.010) (0.010)
Panel B: Average Welfare
183.232 188.221 174.478 165.485
(3.749) (4.072) (4.012) (2.573)
Panel C: Average Welfare for Low-Scored Teachers
176.183 185.154 172.835 156.831
(3.749) (4.072) (9.937) (4.091)
Panel D: Average Welfare for High-Scored Teachers
186.353 188.888 175.951 169.009
(3.251) (3.292) (5.041) (3.653)
Notes: Table reports the welfare metrics for the dynamic behavioral model defined in equations (4.3)–(4.2). Panel A reports structural estimates. The parameter in Panel B (Average Welfare) is defined in Example 2.1. The parameters in Panels C and D (Group Average Welfare) are defined in Example 2.2. Models II and V correspond to the state components in Table 2. Teachers are classified as Low-Scored if their test score is less than or equal to 30, and High-Scored if their score is greater than or equal to 40. The discount factor is $\beta=0.99$. See main text for further details.
5. Dynamic Dual and Doubly Robust Representations of Welfare Metrics
5.1. Dynamic Dual Representation
In this Section, we give a dual representation of the parameter of interest.   This representation is important for several purposes. When $w_{0}(X)$ depends only on $K$, and so is time-invariant, the dual representation gives a simplified formula for $\delta_{0}$ that does not require solving any dynamic problem. Otherwise,  the dual representation leads to a doubly robust moment condition for identification and estimation of the parameter of interest.  The dual representation [^5] See equations (16)-(17) in the first version of the paper Chernozhukov et al. (2019). was derived in the previous version of this paper Chernozhukov, Newey, and Semenova (2019).
A key part of the dual representation is a function of the state variable that is a backward discounted value of $w_{0}(X)$, given by
$\alpha_{0}(X):=\sum_{t\geq 0}\beta^{t}\mathrm{E}[w_{0}(X_{-t})\mid X],$ (5.1)
where $X_{-t}$ is the state variable in period $-t$ in the extended stochastic process ${X_{t}}$ where $t$ ranges over all the integers.  Alternatively, $\alpha_{0}(X)$ is a fixed point of the backward  dynamic operator
[…]
The following result gives the dynamic dual representation of weighted average welfare.
Proposition 5.1 (Dynamic Dual Representation).
Let $V_{\zeta}$ be a net present discounted value of per-period utility $\zeta(x)$ as in (2.1) with $\zeta(\cdot)$ replacing $\zeta_{0}(\cdot)$.  The function $\alpha_{0}(X)$ in equation (5.1) is the unique function such that
$\mathrm{E}[w_{0}(X)V_{\zeta}(X)]=\mathrm{E}[\alpha_{0}(X)\zeta(X)]\quad$ (5.3)
for any $\zeta(X)$ with finite second moment.
An interesting implication of this dual representation is that if $w_{0}(X)$ is time-invariant, then $\delta_{0}$ depends only on the per-period expected utility $\zeta_{0}(X)$.
Corollary 5.1 (Dynamic Dual Representation With Time-Invariant Weight).
If $m(Z,V)=w_{0}(X)V(X)$ and $w_{0}(X)=w_{0}(K)$ depends only on a time-invariant variable $K$ then $\alpha_{0}(X)=(1-\beta)^{-1}w_{0}(K)$ and
$\delta_{0}=\mathrm{E}[w_{0}(X)V_{\zeta}(X)]=(1-\beta)^{-1}\mathrm{E}[w_{0}(K)\zeta(X)].$ (5.4)
In each of our first three examples, the weight was time-invariant so that Corollary 5.1 applies and the parameter of interest $\delta_{0}$ depends only on $\zeta_{0}(X)$. Here are expressions for $\delta_{0}$ for Examples 2.1-2.3.
Example 5.1 (continues=ex:average).
The average welfare is given by
$\delta_{0}=\mathrm{E}[V_{0}(X)]=(1-\beta)^{-1}\mathrm{E}[\zeta_{0}(X)]$ (5.5)
Example 5.2 (continues=ex:group).
The group average welfare is given by
$\delta_{0}=(1-\beta)^{-1}\mathrm{E}[1(K=k)\zeta_{0}(X)]/{\mathrm{P}}(K=k).$ (5.6)
Example 5.3 (continues=ex:stock).
The policy effect is given by
$\delta_{0}=(1-\beta)^{-1}\mathrm{E}[w_{0}(K)\zeta_{0}(X)],w_{0}(K)=[\pi^{*}(K)-\pi(K)]/\pi(K).$ (5.7)
When the weight is not time-invariant, $\alpha_{0}$ will not generally have a closed form or explicit expression because it depends in a complicated way on the dynamic distribution of the state vector $X_{t}$. To help understand better the nature of $\alpha_{0}$ we revisit Example 2.4 where the state variable follows an autoregressive process of order 1 with a Gaussian innovation. While this example may not correspond to a state distribution under a dynamic discrete choice model, we include it for pedagogic purposes to help explain the nature of $\alpha_{0}$.
Proposition 5.2 (Dynamic Dual Representation for Average Marginal Effects).
Consider an AR(1) model with a Gaussian innovation
$S_{t+1}= \rho(K)S_{t}+U_{t},\quad U_{t}\sim\text{IID}N(0,1),$ (5.8)
[…]
$w_{0}(X)=\gamma_{1}(K)S^{2}+\gamma_{0}(K)$ (5.9)
whose intercept $\gamma_{0}(K)$ and the slope $\gamma_{1}(K)$ are functions of the time-invariant type $K$ given in (B.8).   The dynamic dual representation is
[…]
Corollary 5.2.
Consider a white noise model with i.i.d states $S_{t}$, which is a special case of  (5.8) with  $\rho(K)=0$. Then the dynamic dual representation  is time-invariant
[…]
5.2. Doubly Robust Representation
In this section, we give an identifying moment condition for the parameter of interest that is doubly robust in the sense that it holds if just one of $V(\cdot)$ or $\alpha(\cdot)$ is the true function.  This moment condition uses the identifying conditional moment restriction for $V_{0}$ in equation (2.2). Let $Z$ denote a data observation which includes $(X,X_{+})$, $V$ denote a possible value function, and $\lambda(Z,V):=\beta V(X_{+})-V(X)+\zeta_{0}(X)$. Equation (2.2) is equivalent to the conditional moment restriction
$\mathrm{E}[\lambda(Z,V_{0})\mid X]=0$ UNKNOWN (5.11)
This is a nonparametric conditional moment restriction like those of Newey and Powell (2003) and Ai and Chen (2003) where $X_{+}$ is an ”endogenous” variable, $X$ is an ”instrument”, and  $\lambda(Z,V)$ is a nonparametric residual as considered in Chernozhukov et al. (2019) and Chen and Qi (2022). Here we take $\zeta_{0}(X)$ to be a known function and will consider estimation of $\zeta_{0}(X)$ in the next Section.
Let $\alpha$ denote a possible function $\alpha_{0}$. A doubly robust moment function can then be formed as
$g(Z,V,\alpha,\delta)=m(Z,V)-\delta+\alpha(X)\lambda(Z,V).$ (5.12)
Given a function $\xi$ of $X$ define
$\|\xi\|=(\mathrm{E}[\xi(X)^{2}])^{1/2}.$ (5.13)
Lemma 5.1 (Double Robustness of Moment Function  (5.12)).
The moment function satisfies
$\mathrm{E}[g(Z,V,\alpha,\delta_{0})]=\mathrm{E}[(\alpha(X)-\alpha_{0}(X))(\lambda(Z,V)-\lambda(Z,V_{0}))].$ (5.14)
Also
$|\mathrm{E}[g(Z,V,\alpha,\delta_{0})]|\leq(1+\beta)\|\alpha-\alpha_{0}\|\|V-V_{0}\|.$ (5.15)
Lemma 5.1 establishes double robustness of the moment function $g(Z,V,\alpha,\delta)$ which has zero expectation at $\delta=\delta_{0}$ if either $V=V_{0}$ or $\alpha=\alpha_{0}$ by equation  (5.14). A doubly robust estimator of the average treatment effect was given in (Robins and Rotnitzky (1995)) and (Chernozhukov et al. (2022)) characterize doubly robust moment functions as being linear in both non-parameric components.
Example 5.4 (continues=ex:avder).
The doubly robust representation for the average derivative is
[…]
where the true value of $\alpha$ is the backward discounted value (5.1) based on $w_{0}(X)=-\partial_{k}\ln f(K|S)$.
6. Overview of Estimation and Inference
In this Section, we give estimators of the welfare metrics we introduced in Sections 2 and 3. These estimators will account for the estimation of $\zeta_{0}(\cdot)$ and of $w_{0}(\cdot)$ or $m(Z,V)$ by including influence functions for their effect on identifying moments. The inclusion of these influence functions debiases for model selection and/or regularization in the estimation of unknown functions and corrects resulting standard errors for their estimation, as in Chernozhukov et al. (2022).
For simplicity of exposition, we focus on panels with $T=2$ time periods where the pairs of consequent states $(X_{i1},X_{i2})_{i=1}^{n}$ are i.i.d. We use standard cross-fitting for i.i.d data (Schick (1986)) as common in work on debiased machine learning, Chernozhukov et al. (2018). For a weakly dependent time series with $T\geq 3$ periods, cross-fitting along both unit and time dimension is possible by leaving out neighboring folds, as discussed in Semenova et al. (2017).  Related work on conditional moment restrictions with weak dependence includes Chen et al. (2014); Chen and Liao (2015); Chen et al. (2024).
6.1. Average Welfare and Related Averages
We will first consider a weighted average value function parameter with time-invariant weight that is possibly estimated. This case includes average welfare and related averages in Examples 2.12.3. Let $F$ denote an unrestricted distribution for $Z$ and $w(K,F)$ and $\zeta(X,F)$ denote the probability limit (plim) of an estimated weight $\widehat{w}(K)$ and an estimator $\widehat{\zeta}(X)$ respectively. Let $\phi_{w}(Z)$ and  $\phi_{\zeta}(Z)$ be the influence functions of $(1-\beta)^{-1}\mathrm{E}[w(K,F)\zeta_{0}(X)]$ and $(1-\beta)^{-1}E[w_{0}(K)\zeta(X,F)]$ respectively. To nonparametrically debias for the estimation of $w(K)$ and $\zeta(X)$ and so construct a Neyman orthogonal moment function we add $\phi_{w}(Z)$ and $\phi_{\zeta}(Z)$ to the identifying moment function, as in Chernozhukov et al. (2022), to obtain
[…]
$\gamma$ $=(w,\zeta),\phi=(\phi_{w},\phi_{\zeta})$ (6.2)
where the true parameter $\delta_{0}$ solves
$\mathrm{E}[\psi(Z,\gamma_{0},\phi_{0},\delta_{0})]=0$ (6.3)
at the true value $\gamma_{0}$ of $\gamma$ and $\phi_{0}$ of $\phi$.
For cross-fitting purposes, we partition the set of data indices ${1,\ldots,n}$ into $L$ disjoint subsets $I_{\ell}$ of about equal size, $\ell=1,\ldots,L$.   Let $\widehat{\gamma}_{\ell}=(\widehat{w}_{\ell},\widehat{\zeta}_{\ell})$ and $\widehat{\phi}_{\ell}=(\widehat{\phi}_{w\ell},\widehat{\phi}_{\zeta\ell})$ be estimators of the weight, per-period utility, and influence functions constructed using all observations not in  $I_{\ell}$. Also let $\psi(Z,\widehat{\gamma}_{\ell},\widehat{\phi}_{\ell},\delta)$ be as in equation (6.1) with $\widehat{\gamma}_{\ell}$ and $\widehat{\phi}_{\ell}$ in place of $\gamma$ and $\phi$. A cross-fit estimator of $\delta_{0}$ can be obtained from solving  $\sum_{\ell=1}^{L}\sum_{i\in I_{\ell}}\psi(Z_{i},\widehat{\gamma}_{\ell},\widehat{\phi}_{\ell},\delta)/n=0$ for $\delta$ giving
[…]
$\widehat{\Omega}$ $=\frac{1}{n}\sum_{\ell=1}^{L}\sum_{i\in I_{\ell}}\psi^{2}(Z_{i},\widehat{\gamma}_{\ell},\widehat{\phi}_{\ell},\widehat{\delta}).$ (6.5)
An example of estimated per-period utility $\zeta$ and its correction term for dynamic binary choice is given in Appendix A. The Lemma C.1 in Appendix C provides sufficient conditions for the validity of asymptotic inference.
Example 6.1 (continues=ex:average).
The estimate of the average welfare is
$\widehat{\delta}$ $=\frac{1}{n}\sum_{\ell=1}^{L}\sum_{i\in I_{\ell}}[(1-\beta)^{-1}\widehat{\zeta}_{\ell}(X_{i})+\widehat{\phi}_{\zeta\ell}(Z_{i})]$
We next continue with the description of group average welfare in Example 2.2.  A key difference from Example 2.1 is that the time-invariant weighting function $w$  depends on the  group probability ${\mathrm{P}}(K=k)$ which needs to be estimated.
Example 6.2 (continues=ex:group).
Let $\phi_{\zeta}$ be the influence function of $\mathrm{E}[(1-\beta)^{-1}1\{K=k\}\zeta(X,F)/{\mathrm{P}}(K=k)]$.  The weighting function $w(K)=1\{K=k\}/{\mathrm{P}}(K=k)$. The influence function for ${\mathrm{P}}(K=k)$ is
[…]
Thus, the estimator of $\delta_{0}$ reduces to
[…]
whose standard error in (6.5) accounts for estimation of ${\mathrm{P}}(K=k)$ by including the term $\phi_{w}(\cdot)$.
6.2. General case.
In this Section we descrite the estimator of $\delta_{0}=\mathrm{E}[m(Z,V_{0})]$ when $w_{0}(\cdot)$ varies with time. Let $m(Z,V,F)$ denote the plim of the estimated $m(Z,V)$ function, $\phi_{m}(Z)$ the influence function of $\mathrm{E}[m(Z,V_{0},F)]$, and $\phi_{\zeta}(Z)$ the influence function of $\mathrm{E}[\alpha_{0}(X)\zeta(X,F)]$. The orthogonal moment function is
$\psi(Z,\gamma,\phi,\delta)$ $=m(Z,V)-\delta+\alpha(X)(\beta V(X_{+})-V(X)+\zeta(X))+\phi_{m}(Z)+\phi_{\zeta}(Z),$ (6.6)
$\gamma$ $=(m,V,\zeta),\phi=(\alpha,\phi_{m},\phi_{\zeta})$ (6.7)
where $\phi_{m}$ corrects for the estimation of $m(\cdot)$ and $\phi_{\zeta}$ corrects for the estimation of $\zeta$. Algorithm 1 below gives the proposed estimator of the parameter of interest. Lemma C.2 in Appendix C establishes the validity of asymptotic inference.
Example 6.3 (continues=ex:avder).
The estimate of the average derivative is
$\widehat{\delta}=\frac{1}{n}\sum_{\ell=1}^{L}\sum_{i\in I_{\ell}}[\partial_{k}\widehat{V}_{\ell}(X_{i})+\widehat{\alpha}_{\ell}(X_{i})(\beta\widehat{V}_{\ell}(X_{+i})-\widehat{V}_{\ell}(X_{i})+\widehat{\zeta}_{\ell}(X_{i}))+\widehat{\phi}_{\zeta\ell}(Z_{i})].$ (6.8)
Here there is no correction $\phi_{m}$ since the functional $m(Z,V)=\partial_{K}V(X)$ of $V$ does not involve any unknown components.
Example 6.4 (Counterfactual Welfare).
The counterfactual welfare $\delta_{\langle 1\mid 0\rangle}$ in equation (3.6) is a special case of (2.9) with $m(z,V)$ in (3.8) and $w_{0}$ in (3.9). Let $\phi_{\zeta}(Z)$ be the first step influence function (FSIF) of $\mathrm{E}[\alpha_{0}(X)\zeta(X,F)]$ where $\alpha_{0}$ is given in (5.1) based on the weighting function $w_{0}$ in (3.9). The influence function for ${\mathrm{P}}(K=0)$ is
$\phi_{w}(Z)=-\dfrac{\delta_{\langle 1\mid 0\rangle}}{{\mathrm{P}}(K=0)}\left(1\{K=0\}-{\mathrm{P}}(K=0)\right).$
Thus the estimator of $\delta_{1\mid 0}$ reduces to
$\widehat{\delta}_{\langle 1\mid 0\rangle}=\frac{1}{n}\sum_{\ell=1}^{L}\sum_{i\in I_{\ell}}\dfrac{1\{K_{i}=0\}}{\widehat{\mathrm{P}}(K=0)}\widehat{V}_{\ell}(S_{i},1)+\widehat{\alpha}_{\ell}(X_{i})\lambda(Z_{i},\widehat{V}_{\ell})+\widehat{\phi}_{\zeta\ell}(Z_{i})+\widehat{\phi}_{w\ell}(Z_{i}),$
whose standard error in (6.5) accounts for estimation of ${\mathrm{P}}(K=0)$ by including the term $\phi_{w}(\cdot)$.
Algorithm 1 summarizes the estimation steps of welfare metrics.
1:  Partition the set of data indices $\{1,2,\dots,n\}$ into $L$ disjoint subsets of about equal size with where $L$ is an odd number $L\geq 3$. 2:  Estimate the value function by minimizing sample cross-fit criterion function $\widehat{V}_{\ell}=\arg\min_{V}L^{V}_{n}(V)$ 3:  Estimate the dynamic dual representation by minimizing sample cross-fit criterion function $\widehat{\alpha}_{\ell}=\arg\min_{\alpha}L^{\alpha}_{n}(\alpha)$ 4:  Estimate the per-period reward $\zeta(\cdot)$ and correction terms $\widehat{\phi}_{m\ell},\widehat{\phi}_{\zeta\ell}$ using all observations in $I^{c}_{\ell}$ 5:  Estimate the target welfare metric $\delta_{0}$ $\widehat{\delta}$ $=\frac{1}{n}\sum_{\ell=1}^{L}\sum_{i\in I_{\ell}}[m(Z_{i},\widehat{V}_{\ell})+\widehat{\alpha}_{\ell}(X_{i})(\beta\widehat{V}_{\ell}(X_{+i})-\widehat{V}_{\ell}(X_{i})+\widehat{\zeta}_{\ell}(X_{i}))+\widehat{\phi}_{m\ell}(Z_{i})+\widehat{\phi}_{\zeta\ell}(Z_{i})],$ and its standard error $\sqrt{\widehat{\Omega}/n}$ with $\psi$ given in (6.6) as $\widehat{\Omega}$ $=\frac{1}{n}\sum_{\ell=1}^{L}\sum_{i\in I_{\ell}}\psi^{2}(Z_{i},\widehat{\gamma}_{\ell},\widehat{\phi}_{\ell},\widehat{\delta}).$
Algorithm 1 Step 1 partitions the data into $L$ folds to enable sample splitting. Steps 2 and 3 estimate the value function and the dynamic dual representation $\alpha$ on each fold by minimizing cross-fit loss functions. Step 4 estimates the per-period utility $\zeta(\cdot)$ and the bias-correction terms using observations not in fold $\ell$. Step 5 computes the welfare estimate $\widehat{\delta}$ as a bias-corrected plug-in estimator. Step 6 computes the standard error of $\widehat{\delta}$ using the empirical variance of the influence function $\psi$, defined in equation (6.6).
7. Estimation of Value Function and Dynamic Dual Representation
This section introduces novel least squares estimators for both the value function and the dynamic dual representation. In contrast to the welfare metrics, defined in Section 2, these objects do not require strict stationarity to be well-defined. Thus, the estimators of the value function and the dynamic dual representation delivered here do not require the time series to be strictly stationary. Consequently, these estimators apply to any fixed point of a second-kind integral operator.
Sections 7.1 and 7.2 develop a new least squares criterion for the value function that accommodates high-dimensional covariates through penalization, enabling consistent estimation in a high-dimensional state space. Section 7.3 proposes a distinct least squares criterion for the dynamic dual representation, which depends only on the welfare metric of interest. This innovation permits automatic debiasing in the style of Chernozhukov et al. (2022) and Chernozhukov et al. (2024). Section 7.4 discusses the results.
7.1. Least Squares Criterion for Value Function
The starting point of our analysis is the expectation operator $\mathrm{A}_{0}$ defined as
$(\mathrm{A}_{0}\phi)(x):=\beta\mathrm{E}[\phi(X_{+})\mid X=x]$ (7.1)
Rewriting (2.1) in terms of $\mathrm{A}_{0}$ gives
$(\mathrm{I}-\mathrm{A}_{0})V_{0}=\zeta_{0},$ (7.2)
or, equivalently, $V_{0}=(\mathrm{I}-\mathrm{A}_{0})^{-1}\zeta_{0}$. The operator $\mathrm{A}_{0}$ is akin to the integral equation operator in Srisuma and Linton (2012).
The value function can be represented as a minimizer of a criterion function that depends on $\mathrm{A}_{0}$. $V_{0}$ will minimize the expected squared difference of the left and right-hand sides of equation (7.2), that is
$V_{0}$ $=\arg\min_{V}\mathrm{E}[(((\mathrm{I}-\mathrm{A}_{0})V)(X)-\zeta_{0}(X))^{2}]$ (7.3)
$=\arg\min_{V}\mathrm{E}[((\mathrm{I}-\mathrm{A}_{0})V)(X)^{2}-2((\mathrm{I}-\mathrm{A}_{0})V)(X)\zeta_{0}(X)]$ (7.4)
$=\arg\min_{V}\mathrm{E}[((\mathrm{I}-\mathrm{A}_{0})V)(X)^{2}-2(V(X)-\beta V(X_{+}))(X)\zeta_{0}(X)]$ (7.5)
$=\arg\min_{V}\mathrm{E}[(V(X)-\beta V(X_{+}))((\mathrm{I}-\mathrm{A}_{0})V)(X)-2\zeta_{0}(X))],$ (7.6)
where the second equality follows by squaring and dropping the term that does not depend on $V$ and the third and fourth equalities by iterated expectations. The expression minimized following the first equality is the nonparametric two-stage least squares criterion for Newey and Powell (2003), Newey (1991), and Ai and Chen (2003). The expression following the third equality is a hybrid that uses iterated expectations to remove the conditional expectation $\mathrm{A}_{0}$ from all but one term.
Proposition 7.1 (Least Squares Criterion for Value Function).
The value function $V_{0}$ in (2.1) is the unique minimizer of least squares criterion functions
$V_{0}$ $=\arg\min_{V}\mathrm{E}[((\mathrm{I}-\mathrm{A}_{0})V)(X)^{2}-2(V(X)-\beta V(X_{+}))(X)\zeta_{0}(X)]$ (7.7)
$=\arg\min_{V}\mathrm{E}[(V(X)-\beta V(X_{+}))((\mathrm{I}-\mathrm{A}_{0})V)(X)-2\zeta_{0}(X))].$ (7.8)
Proposition 7.1 gives two least squares criterion functions. We use (7.7) and (7.8) to construct a Lasso and a Neural Network estimator, respectively. We describe the Lasso estimator in Section 7.2 and Neural Network estimator in Appendix D.
To describe the Lasso estimator of the value function let
$b(x)=(b_{1}(x),\dots,b_{p}(x))\in\mathrm{R}^{p},$
be a vector of basis functions. We approximate the value function using a linear form
$V(x)\approx\sum_{j=1}^{p}b_{j}(x)\rho_{Vj}=b(x)^{\prime}\rho_{V},$
where $\rho_{V}=(\rho_{V1},\rho_{V2},\dots,\rho_{Vp})\in\mathrm{R}^{p}$ is a $p$-vector of coefficients. The vector is chosen to minimize an approximate least squares criterion
$\rho_{V}=\arg\min_{\rho\in\mathrm{R}^{p}}\rho^{\prime}G^{V}\rho-2M^{V}\rho$
where $G^{V}$ is a symmetric $p\times p$ matrix
$G^{V}$ $=\mathrm{E}[((\mathrm{I}-\mathrm{A}_{0})b)(X)((\mathrm{I}-\mathrm{A}_{0})b)^{\prime}(X)]$ (7.9)
and $M^{V}$ is a linear term
$M^{V}$ $=\mathrm{E}[(b(X)-\beta b(X_{+}))\zeta_{0}(X)].$
The FOC reduces to
$G^{V}\rho_{V}=M^{V}.$
We choose the criterion (7.7) as opposed to (7.8) so that the sample version of matrix $G^{V}$ is symmetric and positive-definite.
Given an i.i.d sample $(X_{i},X_{i+})_{i=1}^{n}$, we construct a sample estimate of $\rho_{V}$ in the regime where $\dim(\rho_{V})=p_{V}\gg n$. For simplicity of exposition, we abstract away from subsequent estimation steps and drop respective cross-fitting indices. Given a plug-in estimate $\widehat{\mathrm{A}}b$ of $\mathrm{A}_{0}b$ and $\widehat{\zeta}$ and $\zeta_{0}$ estimated on a hold-out sample, define
$\widehat{G}^{V}$ $=n^{-1}\sum_{i=1}^{n}(b(X_{i})-(\widehat{\mathrm{A}}b)(X_{i}))(b(X_{i})-(\widehat{\mathrm{A}}b)(X_{i}))^{\prime},$
$\widehat{M}^{V}$ $=n^{-1}\sum_{i=1}^{n}(b(X_{i})-\beta b(X_{i+}))\widehat{\zeta}(X_{i})$
Given a radius $\rho_{V}$, an $\ell_{1}$-regularized estimator of the value function takes the form
$\widehat{V}(x)$ $=b(x)^{\prime}\widehat{\rho}_{V},$ (7.10)
$\widehat{\rho}_{V}$ $=\arg\min_{\rho\in\mathrm{R}^{p}}\rho^{\prime}\widehat{G}^{V}\rho-2\widehat{M}^{V}\rho+r_{V}\|\rho\|_{1}.$ (7.11)
7.2. Mean Square Convergence for Lasso Estimator
Assumption 7.1 requires that $V_{0}$ belongs to the mean square closure $\Gamma$ of linear combinations $b(x)^{\prime}\bar{\rho}$, as well as that the approximating coefficients $\bar{\rho}$ are sufficiently sparse.
Assumption 7.1.
(1) There exist constants $C>1$ and $\xi_{V}>1/2$ and $d_{V}\in(0,1/2)$ such that for each positive integer $s_{V}\leq Cn^{-2(d_{V})/(2\xi_{V}+1)}$ there is $\bar{\rho}$ with $s_{V}$ nonzero elements such that
$\mathrm{E}[(V_{0}(X)-b^{\prime}(X)\bar{\rho})^{2}]\leq Cs_{V}^{-2\xi_{V}}.$
(2) The matrix $\mathrm{E}[b(X)b(X)^{\prime}]$ is positive definite with eigenvalues bounded from above by $\bar{\lambda}$ and below by $\underline{\lambda}>0$. (3) $\sup_{j}|b_{j}(X)|$ are bounded a.s. (4) The radius $r_{V}$ is chosen such that $\varepsilon_{n}=o(r_{V}),$ $r_{V}=o(n^{c}\varepsilon_{n})$ for all $c>0$, and there exists $C>0$ such that $p_{V}\leq Cn^{C}.$ (5) The first-stage estimators $\widehat{\zeta}$ of $\zeta_{0}$ and $\widehat{\mathrm{A}}b$ of $\mathrm{A}_{0}b$ converge as $\|\widehat{\zeta}-\zeta_{0}\|=o_{P}(\zeta_{n})$ and $\|\widehat{\mathrm{A}}-\mathrm{A}_{0}\|=o_{P}(a_{n})$ where $\zeta_{n}+a_{n}=o(n^{-d_{V}})$ for some positive constant $d_{V}\in(0,1/2)$.
Assumption 7.1(1) requires value function to be approximately sparse in the chosen basis. Assumptions 7.1(2)-(4) are standard regularity conditions. Assumption 7.1(5) reduces to a rate condition on the first-stage estimators.
Theorem 7.1 (Mean Square Rate for Lasso estimator of Value Function).
If Assumption 7.1 holds, then, for any $c>0$,
$\|\widehat{V}-V_{0}\|=o_{p}(n^{c}n^{-2d_{V}\xi_{V}/(2\xi_{V}+1)}).$ (7.12)
Theorem 7.1 gives a mean square convergence rate for the value function. The rate is determined by the sparsity parameter $\xi_{V}$ of the value function and the first-stage rate parameter $d_{V}\in(0,1/2)$.
Remark 7.1 (Verification of Rate Condition (5) in Assumption 7.1).
To give an example of operator $\mathrm{A}_{0}$, we consider the dynamic discrete choice problem discussed in Section 2 where the state transition is deterministic conditional on action. Then the estimator of $\mathrm{A}_{0}$ reduces to an estimator of choice probabilities, and the rate $a_{n}=O(p_{n})$ where $p_{n}$ is the $\ell_{\infty}$-rate for CCPs.
Remark 7.2 (Lasso estimator of $\mathrm{A}_{0}$).
The following example of $\mathrm{A}_{0}$ is based on a Lasso estimator. For $j=1,2,\dots,p$, define the estimators as
$\widehat{G}$ $=n^{-1}\sum_{i=1}^{n}b(X_{i})b^{\prime}(X_{i}),\quad\widehat{M}_{j}=n^{-1}\sum_{i=1}^{n}b(X_{i})b_{j}(X_{+i}).$ (7.13)
$\widehat{\rho}_{j}$ $=\arg\min_{\rho}\rho^{\prime}\widehat{G}\rho-2\widehat{M}_{j}\rho+r_{j}\|\rho\|_{1}$ (7.14)
where
$(\widehat{\mathrm{A}}b)_{j}=b(x)^{\prime}\widehat{\rho}_{j},\quad j=1,2,\dots,p.$
Then, the rate condition (5) in Assumption 7.1 on $a_{n}$ can be verified by the first-case rate of $\sup_{1\leq j\leq p}\|\widehat{\rho}_{j}-\rho_{j}\|$ and can be established using the tools of Semenova et al. (2017).
7.3. Least Squares Criterion for Dynamic Dual Representation
In this Section, we derive dynamic dual criterion function. Define
$(\mathrm{A}_{0}^{*}\phi)(x):=\beta\mathrm{E}[\phi(X_{-})\mid X=x]$ (7.15)
From the dual representation of (5.2) we know that $\alpha_{0}$ satisfies
$(\mathrm{I}-\mathrm{A}_{0}^{*})\alpha_{0}=w_{0}$ (7.16)
and $\mathrm{I}-\mathrm{A}^{*}_{0}$ is invertible.
We show that $\alpha_{0}$ can be represented as a minimizer of a criterion function that depends on $\mathrm{A}^{*}_{0}$. $\alpha_{0}$ will minimize the expected squared difference of the left and right-hand sides of equation (7.16), that is
$\alpha_{0}$ $=\arg\min_{\alpha}\mathrm{E}[(((\mathrm{I}-\mathrm{A}^{*}_{0})\alpha)(X)-w_{0}(X))^{2}]$ (7.17)
$=\arg\min_{\alpha}\mathrm{E}[((\mathrm{I}-\mathrm{A}^{*}_{0})\alpha)(X)^{2}-2((\mathrm{I}-\mathrm{A}^{*}_{0})\alpha)(X)w_{0}(X)]$
$=\arg\min_{\alpha}\mathrm{E}[((\mathrm{I}-\mathrm{A}^{*}_{0})\alpha)(X)^{2}-2m(Z,(\mathrm{I}-\mathrm{A}^{*}_{0})\alpha)]$ (7.18)
$=\arg\min_{\alpha}\mathrm{E}[(\alpha(X)-\beta\alpha(X_{-}))((\mathrm{I}-\mathrm{A}^{*}_{0})\alpha)(X)-2m(Z,(\mathrm{I}-\mathrm{A}^{*}_{0})\alpha)],$ (7.19)
where the second equality follows by squaring and dropping the term that does not depend on $\alpha$, the third equality follows the Riesz representation in (2.11), and the third equality by iterated expectations.
Proposition 7.2 (Least Squares Criterion for Dynamic Dual Representation).
The dynamic dual representation $\alpha_{0}$ in (5.2) is the unique minimizer of the quadratic criterion function
$\alpha_{0}$ $=\arg\min_{\alpha}\mathrm{E}[((\mathrm{I}-\mathrm{A}^{*}_{0})\alpha)(X)^{2}-2m(Z,(\mathrm{I}-\mathrm{A}^{*}_{0})\alpha)]$ (7.20)
$=\arg\min_{\alpha}\mathrm{E}[((\mathrm{I}-\mathrm{A}^{*}_{0})\alpha)((\mathrm{I}-\mathrm{A}^{*}_{0})\alpha)(X)-2m(Z,(\mathrm{I}-\mathrm{A}^{*}_{0})\alpha)].$ (7.21)
$=\arg\min_{\alpha}\mathrm{E}[(\alpha(X)-\beta\alpha(X_{-}))((\mathrm{I}-\mathrm{A}^{*}_{0})\alpha)(X)-2m(Z,(\mathrm{I}-\mathrm{A}^{*}_{0})\alpha)].$ (7.22)
Given a vector of basis functions
$b(x)=(b_{1}(x),b_{2}(x),\dots,b_{j}(x),\dots,b_{p}(x))\in\mathrm{R}^{p},$
where $p$ can differ from the one Section 7.1, we approximate the dynamic dual representation via a linear form
$\alpha(x)\approx\sum_{j=1}^{p}b_{j}(x)\rho_{\alpha j}=b(x)^{\prime}\rho_{\alpha}$
where $\rho_{\alpha}=(\rho_{\alpha 1},\rho_{\alpha 1},\dots,\rho_{\alpha p})\in\mathrm{R}^{p}$ is a $p$-vector. The vector is chosen to minimize an approximate least squares criterion
$\rho_{\alpha}=\arg\min_{\rho\in\mathrm{R}^{p}}\rho^{\prime}G_{\alpha}\rho-2M_{\alpha}\rho$
where the $p\times p$ matrix is an outer product of
$G_{\alpha}$ $=\mathrm{E}[((\mathrm{I}-\mathrm{A}^{*}_{0})b)(X)((\mathrm{I}-\mathrm{A}^{*}_{0})b)^{\prime}(X)]$
and the free term is
$M_{\alpha}$ $=\mathrm{E}[m(Z,(\mathrm{I}-\mathrm{A}^{*}_{0})\alpha)]$
We choose the criterion (7.21) as opposed to (7.22) so that the sample version of matrix $G_{\alpha}$ is symmetric and positive-definite. Given a radius $r_{\alpha}$, an $\ell_{1}$-regularized minimum distance estimator of the dynamic dual representation
$\widehat{\alpha}(x)$ $=b(x)^{\prime}\widehat{\rho}_{\alpha}$ (7.23)
$\widehat{\rho}_{\alpha}$ $=\arg\min_{\rho\in\mathrm{R}^{p}}\rho^{\prime}\widehat{G}^{\alpha}\rho-2\widehat{M}^{\alpha}\rho+r_{\alpha}\|\rho\|_{1}$ (7.24)
where
$\widehat{G}^{\alpha}$ $=n^{-1}\sum_{i=1}^{n}(\mathrm{I}-\widehat{\mathrm{A}}^{*})b(X_{i})(\mathrm{I}-\widehat{\mathrm{A}}^{*})b^{\prime}(X_{i}),\qquad\widehat{M}^{\alpha}=n^{-1}\sum_{i=1}^{n}m(Z_{i},(\mathrm{I}-\widehat{\mathrm{A}}^{*})b).$
Theorem 7.2 (Mean Square Rate for Lasso estimator of Dynamic Dual Representation).
Suppose Assumption 7.1 holds with $\alpha_{0}$ in place of $V_{0}$ and $\xi_{\alpha}$ in place of $\xi_{V}$, and suppose $\xi_{\alpha}>1/2$ and $a^{*}_{n}=o(n^{-d_{\alpha}})$. Then
$\|\widehat{\alpha}-\alpha_{0}\|=o_{p}(n^{c}n^{(-2d_{\alpha}\xi_{\alpha})/(2\xi_{\alpha}+1)}).$ (7.25)
Theorem 7.2 gives a mean square convergence rate for dynamic dual representation. The rate is determined by the sparsity parameter $\xi_{\alpha}$ of the dynamic dual representation and the first-stage rate parameter $d_{\alpha}$. The estimator is automatic in the sense that it only requires knowledge of the linear functional $m(Z,(\mathrm{I}-\mathrm{A}^{*})b)$.
7.4. Discussion and Related Results
In this Section, we discuss related results. Remark 7.3 discusses cross-fitting. Remark 7.4 introduces a neural network estimator of the value function. Remark 7.5 introduces a neural network estimator of the dynamic dual representation. Remark 7.6 describes the automatic property of the dynamic dual criterion function. Remark 7.7 verifies rate conditions for asymptotic theory. Remark 7.8 sketches a Dantzig estimator of the value function.
Remark 7.3 (Cross-fitting).
To ensure that the nuisance components $\zeta_{0},\mathrm{A}_{0},\mathrm{A}^{*}_{0}$ and their respective criterion functions are estimated on different samples, the standard cross-fitting procedure is modified as follows. Let $L\geq 3$ be the number of partitions. For each partition $\ell\in\{1,2,\dots,L\}$, let $I^{c}_{\ell}=(Z_{i})_{i\notin I_{\ell}}$ denote the set observations not in $I_{\ell}$. Partition $I^{c}_{\ell}=I^{c1}_{\ell}\sqcup I^{c2}_{\ell}$ into two halves. For each nuisance parameter $\gamma\in\{\zeta,\mathrm{A},\mathrm{A}^{*}\}$, let $\widehat{\gamma}^{1}_{\ell},\gamma^{2}_{\ell}$ denote the estimator computed on $I^{c1}_{\ell}$ and $I^{c2}_{\ell}$, respectively. A cross-fit criterion for the value function is
$L^{V}_{n}(V)=\sum_{i\in I^{c1}_{\ell}}\ell_{V}(Z_{i},V,\widehat{\omega}^{2}_{\ell})+\sum_{i\in I^{c2}_{\ell}}\ell_{V}(Z_{i},V,\widehat{\omega}^{1}_{\ell}),\quad\omega=(\zeta,\mathrm{A}).$
In a special case of Lasso estimator, the criterion $L^{V}_{n}(V)$ reduces to
$L^{V}_{n}(\rho)$ $=\rho^{\prime}\widehat{G}^{V}\rho-2\widehat{M}^{V}\rho+r_{V}\|\rho\|_{1}$
where
$\widehat{G}^{V}$ $=\sum_{i\in I^{c1}_{\ell}}(\mathrm{I}-\widehat{\mathrm{A}}^{2}_{\ell})b(X_{i})(\mathrm{I}-\widehat{\mathrm{A}}^{2}_{\ell})b^{\prime}(X_{i})+\sum_{i\in I^{c2}_{\ell}}(\mathrm{I}-\widehat{\mathrm{A}}^{1}_{\ell})b(X_{i})(\mathrm{I}-\widehat{\mathrm{A}}^{1}_{\ell})b^{\prime}(X_{i})$
$\widehat{M}^{V}$ $=\sum_{i\in I^{c1}_{\ell}}(b(X_{i})-\beta b(X_{i+}))\widehat{\zeta}^{2}_{\ell}(X_{i})+\sum_{i\in I^{c2}_{\ell}}(b(X_{i})-\beta b(X_{i+}))\widehat{\zeta}^{1}_{\ell}(X_{i}).$
Remark 7.4 (Neural network estimator of value function).
Proposition 7.1 facilitates a general plug-in estimator of the value function with an arbitrary function class. Given a first-stage estimator of the per-period utility $\widehat{\zeta}$ of $\zeta_{0}$ and the expectation operator $\widehat{\mathrm{A}}$ of $\mathrm{A}_{0}$, define
$\widehat{V}:=\arg\min_{V\in\mathcal{V}_{n}}n^{-1}\sum_{i=1}^{n}\ell_{V}(Z_{i},V,\widehat{\zeta},\widehat{\mathrm{A}}),$ (7.26)
where $\ell_{V}(Z,V,\zeta_{0},\mathrm{A}_{0})$ is taken to as in (7.6). Theorem E.1 in Appendix E establishes mean square consistency of $\widehat{V}$ to $V_{0}$ for an arbitrary function class. Corollary E.1 establishes mean square consistency of the neural network estimator of $V_{0}$.
Remark 7.5 (Neural network estimator of dynamic dual representation).
Proposition 7.2 facilitates a general plug-in estimator of dynamic dual representation with an arbitrary function class. Given a first-stage estimator of the operator $\widehat{\mathrm{A}}^{*}$ of $\mathrm{A}^{*}_{0}$, define
$\widehat{\alpha}:=\arg\min_{\alpha\in\mathcal{A}_{n}}n^{-1}\sum_{i=1}^{n}\ell_{\alpha}(Z_{i},\alpha,\widehat{\mathrm{A}}^{*}).$ (7.27)
Theorem E.2 in Appendix E establishes mean square consistency of $\widehat{\alpha}$ to $\alpha_{0}$ for an arbitrary function class. It can be used to derive mean square consistency of neural network estimator of $\alpha_{0}$.
Remark 7.6 (Automatic property of criterion function (7.22)).
The criterion function (7.22) have a convenient property that it only depends on the parameter of interest through $m(Z,(\mathrm{I}-\mathrm{A}_{0})\alpha)$ and does not require an explicit formula for the weighting function $w_{0}$. A similar property has been establishes for the results in Chernozhukov et al. (2022) and Chernozhukov et al. (2024). When the model is static, that is $\beta=0$, Theorem 7.2 recovers Theorem 1 of Chernozhukov et al. (2022). Likewise, the criterion function (7.22) reduces to
$\alpha_{0}=\arg\min_{\alpha}\mathrm{E}[\alpha^{2}(X)-2m(Z,\alpha)]$
proposed in Chernozhukov et al. (2024).
Remark 7.7 (Verification of rate conditions for asymptotic theory.).
Suppose the conditions of Theorems 7.1 and 7.2 hold with $\xi_{V},\xi_{\alpha}$ and $d_{V},d_{\alpha}$. Furthermore, suppose
$22(d_{\alpha}\xi_{\alpha})/(2\xi_{\alpha}+1)+2(d_{V}\xi_{V})/(2\xi_{V}+1)>1/$ 2 .
For example, if $\min(\xi_{\alpha},\xi_{V})>1+0.01$ and $\min(d_{V},d_{\alpha})\in(3/8,1/2)$, the product of mean square rates can be upper bounded as
$n^{1/2}\|\widehat{V}_{\ell}-V_{0}\|\|\alpha_{\ell}-\alpha_{0}\|=o_{P}(n^{1/2}n^{-1/4}\cdot n^{-1/4})=o_{P}(1)$
which suffices for the product rate condition
$n^{1/2}\|\widehat{V}_{\ell}-V_{0}\|\|\alpha_{\ell}-\alpha_{0}\|=o_{P}(1)$
of Assumption C.5 in Appendix C.
Remark 7.8 (Dantzig estimator of the value function).
Earlier work has focused on an instrumental variable (IV) approach corresponding to a minimum distance problem
$\mathrm{E}[b(X)(b(X)-\beta b(X_{+}))^{\prime}]\rho_{V}-\mathrm{E}[b(X)\zeta_{0}(X)]$ (7.28)
$=G\rho_{V}-M=0.$
Unlike $G_{V}$ in (7.9), the matrix $G=\mathrm{E}[b(X)(b(X)-\beta b(X_{+}))^{\prime}]$ is not symmetric and does not correspond to a FOC of a least squares criterion function. Yet, a sparsity restriction of $\rho_{V}$ can be imposed using a Dantzig-type estimator
$\min_{\rho\in\mathrm{R}^{p}}\|\rho\|_{1}\text{ subject to }\|\widehat{G}\rho-\widehat{M}\|_{\infty}\leq\lambda_{n},$ (7.29)
where $\lambda_{n}$ is an appropriately chosen sequence.
8. Conclusions
In this paper we introduce welfare metrics – including welfare decompositions into direct and indirect effects – and give a complete set of estimation and inference results for them in the presence of high-dimensional state space. The results are presented for the dynamic binary choice model of Rust (1987); Hotz and Miller (1993); Aguirregabiria and Mira (2002) but are applicable to other dynamic models (e.g., dynamic games Aguirregabiria and Mira (2007); Bajari et al. (2007)). For the case of average welfare and related metrics, the proposed estimator is a known function of choice probabilities and the structural parameter. In particular, if the model has “terminal action” property, value function or any other dynamic object does not have to be estimated at all. We have applied these methods to estimate the average teachers welfare in an application to teachers absenteeism as in Duflo et al. (2012).
Appendix A Correction Term for Expected Per-Period Utility
In Section 6, we introduce the estimator of welfare metric parameter $\delta_{0}$. In this Section, we derive the correction term $\phi_{\zeta}$ for the expected per-period utility. We focus on the dynamic binary choice model in Section 2. The correction term is a sum of two terms
$\phi_{\zeta}(Z)=\phi_{p}(Z)+\phi_{\theta}(Z)$ (A.1)
where $\phi_{p}(Z)$ accounts for estimation of CCPs and $\phi_{\theta}(Z)$ accounts for estimation of $\theta$. Since $p(x)$ is a conditional mean function, the correction term follows from Newey (1994)
$\phi_{p}(Z)=\alpha(X)(u(X,1)-u(X,0)+\ln(1-p(X))-\ln p(X))(J-p(X)).$ (A.2)
To specify the correction term $\phi_{\theta}(Z)$, we impose the following assumption.
Assumption A.1 (Structural parameter).
Suppose utilities take a linear index form
$u(x,1)=D_{1}(x)^{\prime}\theta_{11},\quad u(x,0)=D_{0}(x)^{\prime}\theta_{10},$ (A.3)
where $\theta_{0}=(\theta_{11},\theta_{10})$. We assume that $\theta_{0}$ is identified via a moment equation
$\mathrm{E}[g_{\theta}(Z,\theta_{0})]=0.$
Furthermore, there exists an stimator $\widehat{\theta}$ of $\theta_{0}$ that is asymptotically linear
$\sqrt{n}(\widehat{\theta}-\theta_{0})=\mathrm{E}_{n}\psi_{\theta}(Z_{i},\theta_{0})+o_{P}(1),$
where $\psi_{\theta}$ is an influence function of $\theta_{0}$.
Assumption A.1 simplifies our exposition. For the case of Rust (1987) model, this assumption is satisfied for the debiased estimator proposed in e.g. Chernozhukov et al. (2022) under some conditions. Then, invoking Chernozhukov et al. (2015) gives the correction term for $\theta$ as
$\phi_{\theta}(Z)=\Gamma_{\theta}^{\prime}\Omega_{\theta}^{-1}g_{\theta}(Z,\theta)$
where the matrices are
$\Gamma_{\theta}$ $=(\mathrm{E}[\alpha(X)D_{1}(X)p(X)],\mathrm{E}[\alpha(X)D_{0}(X)(1-p(X))])^{\prime},\qquad\Omega_{\theta}=\text{Var}(g_{\theta}(Z,\theta_{0})).$
Appendix B Proofs for Results referenced in Main Text
Lemma B.1 (Equivalence).
The fixed point of (2.2) coincides with the net present discounted value (2.1).
Proof of Lemma B.1.
Decomposing value function into the current $(t=0)$ and the future $(t\geq 1)$ periods gives
$V_{0}(X)$ $=\zeta_{0}(X)+\sum_{t=1}^{\infty}\beta^{t}\mathrm{E}[\zeta_{0}(X_{t})\mid X]$
Rearranging time indices in the continuation value and concentrating $\beta$ out
$V_{0}(X)=\zeta_{0}(X)+\beta\sum_{t=0}^{\infty}\beta^{t}\mathrm{E}[\zeta_{0}(X_{t+1})\mid X].$
By Law of Iterated Expectations, the continuation value can be represented as a conditional expectation
$\sum_{t=0}^{\infty}\beta^{t}\mathrm{E}[\zeta_{0}(X_{t+1})\mid X]=\mathrm{E}[\sum_{t=0}^{\infty}\beta^{t}\mathrm{E}[\zeta_{0}(X_{t+1})\mid X_{+}]\mid X]=\mathrm{E}[V_{0}(X_{+})\mid X].$
Putting the summands together gives (2.2). $\blacksquare$
Lemma B.2 (Theorem 4.8, Kress (1989)).
Let $\mathcal{L}_{2}$ be the Hilbert space with the inner product $\langle f,g\rangle=\mathrm{E}[f(X)g(X)]$. Then the following statement holds. (1) The operator $F(\zeta):\mathcal{L}_{2}\rightarrow\mathrm{R}$
$F(\zeta):$ $=\mathrm{E}[w(X)V_{\zeta}(X)]=\mathrm{E}\bigg{[}w(X)\left(\sum_{t\geq 0}\beta^{t}\mathrm{E}[\zeta(X_{t})\mid X]\right)\bigg{]}$ (B.1)
is a linear operator
$F(\alpha\zeta_{1}+\beta\zeta_{2})=\alpha F(\zeta_{1})+\beta F(\zeta_{2}).$ (B.2)
(2) $F(\zeta)$ is a mean square continuous functional of $\zeta$, that is,
$|F(\zeta)|$ $\leq\|w\|(1-\beta)^{-1}\|\zeta\|_{2}\quad\forall\zeta.$ (B.3)
(3) There exists a unique $\alpha_{0}\in\mathcal{L}_{2}$ such that $F(\zeta)=\mathrm{E}[\alpha_{0}(X)\zeta(X)]$ for any $\zeta\in\mathcal{L}_{2}$
Proof of Lemma B.2.
(1) Linearity of (B.2) follows from linearity of expectation. (2): We have that
$\|V_{\zeta}(X)\|$ $=\|\sum_{t\geq 0}\beta^{t}\mathrm{E}[\zeta(X_{t})\mid X]\|.$
Triangular inequality for the norm, the properties of expectation, and stationarity
$\|V_{\zeta}(X)\|$ $\leq\sum_{t\geq 0}\beta^{t}\|\mathrm{E}[\zeta(X_{t})\mid X]\|$
$\leq\sum_{t\geq 0}\beta^{t}\|\zeta\|\leq(1-\beta)^{-1}\|\zeta\|.$
Cauchy Schwarz inequality implies
$|F(\zeta)|$ $\leq\|w\|\|V_{\zeta}\|\leq\|w\|(1-\beta)^{-1}\|\zeta\|.$
Therefore, $F(\zeta)$ is bounded. (3) follows from Riesz representation lemma. $\blacksquare$
Proof of Proposition 3.1.
The counterfactual welfare can be represented as
[…]
Rewriting it in an expectation form is
$\delta_{\langle 1\mid 0\rangle}$ $=\mathrm{E}\bigg{[}V_{0}^{1}(S)\dfrac{\pi^{0}(S)}{\pi^{1}(S)}\mid K=1\bigg{]}=^{(i)}\mathrm{E}\bigg{[}V_{0}(X)\dfrac{\pi^{0}(S)}{\pi^{1}(S)}\mid K=1\bigg{]}$
$=\dfrac{\mathrm{E}\bigg{[}1\{K=1\}V_{0}(X)\dfrac{\pi^{0}(S)}{\pi^{1}(S)}\bigg{]}}{{\mathrm{P}}(K=1)}=^{(ii)}\mathrm{E}[w_{0}(X)V_{0}(X)]$
which coincides with (5.9) with $w_{0}$ in (3.9). $\blacksquare$
Proof of Lemma 5.1.
Let $\zeta$ be any integrable function and $V_{\zeta}(X)$ be defined in (2.1). We note that
$\delta_{0}=\mathrm{E}[w_{0}(X)V_{\zeta}(X)]$ $=^{(i)}\sum_{t=0}^{\infty}\beta^{t}\mathrm{E}[w_{0}(X)\zeta(X_{t})]$
$=^{(ii)}\sum_{t=0}^{\infty}\beta^{t}\mathrm{E}[w_{0}(X_{-t})\zeta(X)]$
$=^{(iii)}\mathrm{E}[\alpha_{0}(X)\zeta(X)],$
where (i) follows from the definition of $V(X)$ in equation (2.1), (ii) from strict stationarity and
$\mathrm{E}[w_{0}(X)\zeta(X_{t})]=\mathrm{E}[w_{0}(X_{-t})\zeta(X)]$
for any $t$ and (iii) from the definition of dynamic Riesz representer $\alpha_{0}(X)$ is in (5.1). Rewriting (5.1) and separating $w_{0}(X)$ from the rest of the terms gives
$\alpha_{0}(X)$ $=w_{0}(X)+\sum_{t\geq 1}\beta^{t}\mathrm{E}[w_{0}(X_{-t})\mid X]$
$=w_{0}(X)+\beta\sum_{t\geq 0}\beta^{t}\mathrm{E}[w_{0}(X_{-(t+1)})\mid X]$
$=w_{0}(X)+\beta\mathrm{E}[\sum_{t\geq 0}\beta^{t}\mathrm{E}[w_{0}(X_{-(t+1)})\mid X_{-1}]\mid X]=w_{0}(X)+\beta\mathrm{E}[\alpha_{0}(X_{-1})\mid X].$
Replacing $X$ by $X_{+}$ gives $\alpha_{0}(X)$ as a solution to an integral equation (5.2). Uniqueness of $\alpha_{0}$ follows from Riesz representation lemma whose sufficient conditions are verified in Lemma B.2.
$\blacksquare$
Proof of Corollary 5.1.
Plugging $w_{0}(X)=w_{0}(K)$ into equation (5.1) for $\alpha_{0}(X)$ gives
$\alpha_{0}(X)=\sum_{t\geq 0}\beta^{t}\mathrm{E}[w_{0}(X_{-t})\mid X]=\sum_{t\geq 0}\beta^{t}\mathrm{E}[w_{0}(K)\mid X]=(1-\beta)^{-1}w_{0}(K).$ (B.4)
$\blacksquare$
Proof of Lemma 5.1.
Let $V_{0}(X)$ be the value function defined in (2.2), and $V(X)$ be any other integrable function. Define the difference between the two
$\Delta_{V}(X):=V(X)-V_{0}(X).$
Since $\lambda(Z,\cdot)$ is linear in $V$,
$\lambda(Z,V)-\lambda(Z,V_{0})$ $=\beta\Delta_{V}(X_{+})-\Delta_{V}(X).$ (B.5)
Step 1 . We define the error terms. Decomposing
$\mathrm{E}[g(Z,V,\alpha,\delta_{0})]=$ $\mathrm{E}[g(Z,V,\alpha,\delta_{0})-g(Z,V_{0},\alpha_{0},\delta_{0})]$
$=\mathrm{E}[m(Z,V)-m(Z,V_{0})]+\mathrm{E}[\alpha(X)\lambda(Z,V)-\alpha_{0}(X)\lambda(Z,V_{0})]$
$=\mathrm{E}[m(Z,V)-m(Z,V_{0})]$
$+\mathrm{E}[\alpha_{0}(X)(\lambda(Z,V)-\lambda(Z,V_{0}))]$
$+\mathrm{E}[(\alpha(X)-\alpha_{0}(X))\lambda(Z,V)].$
Invoking Riesz representation (2.11) gives $\mathrm{E}[m(Z,V)-m(Z,V_{0})]=\mathrm{E}[w_{0}(X)\Delta_{V}(X)]=:S_{1}$. Plugging (B.5) into the second term
$\mathrm{E}[\alpha_{0}(X)(\lambda(Z,V)-\lambda(Z,V_{0}))]=\mathrm{E}[\beta\alpha_{0}(X)\Delta_{V}(X_{+})-\alpha_{0}(X)\Delta_{V}(X)]=:S_{2}+S_{3}.$
Then
$\mathrm{E}[g(Z,V,\alpha,\delta_{0})]=:S_{1}+S_{2}+S_{3}+\mathrm{E}[(\alpha(X)-\alpha_{0}(X))\lambda(Z,V)].$
Step 2 . We show that $S_{1}+S_{2}+S_{3}=0$, which suffices to prove (5.14). Invoking stationarity allows to replace $X$ by $X_{+}$ in the first and third summand
$S_{1}=\mathrm{E}[w(X_{+})\Delta_{V}(X_{+})],\qquad S_{3}=\mathrm{E}[\alpha(X_{+})\Delta_{V}(X_{+})],$
which allows concentrating out $\Delta_{V}(X_{+})$, that is
$S_{1}+S_{2}+S_{3}$ $=\mathrm{E}[(w(X_{+})+\beta\alpha_{0}(X)-\alpha_{0}(X_{+}))\Delta_{V}(X_{+})].$
By definition of $\alpha_{0}(X)$ as a fixed point of (5.2), $S_{1}+S_{2}+S_{3}=0$ which gives (5.14). Since $\mathrm{E}[(\alpha(X)-\alpha_{0}(X))\lambda(Z,V_{0})]=0$, the final term can be written as the product
$\mathrm{E}[(\alpha(X)-\alpha_{0}(X))\lambda(Z,V)]=\mathrm{E}[(\alpha(X)-\alpha_{0}(X))(\lambda(Z,V)-\lambda(Z,V_{0}))],$
which gives (5.14).
Step 3 . We prove (5.15). Invoking Cauchy Schwartz gives
$|\mathrm{E}[(\alpha(X)-\alpha_{0}(X))\Delta_{V}(X)]|\leq\|\alpha-\alpha_{0}\|\|V-V_{0}\|.$
Invoking Cauchy Schwartz and stationarity gives
$|\mathrm{E}[(\alpha(X)-\alpha_{0}(X))\Delta_{V}(X_{+})]|$ $\leq(\mathrm{E}[(\alpha(X)-\alpha_{0}(X))^{2}])^{1/2}(\mathrm{E}[(V(X_{+})-V_{0}(X_{+}))^{2}])^{1/2}$
$\leq\|\alpha-\alpha_{0}\|\|V-V_{0}\|.$
Adding the terms together gives an upper bound
$|\mathrm{E}[g(Z,V,\alpha,\delta_{0})]|$ $\leq\beta|\mathrm{E}[(\alpha(X)-\alpha_{0}(X))\Delta_{V}(X_{+})]|+|\mathrm{E}[(\alpha(X)-\alpha_{0}(X))\Delta_{V}(X)]|$
$\leq(\beta+1)\|\alpha-\alpha_{0}\|\|V-V_{0}\|,$
which coincides with (5.15).
$\blacksquare$
Proof of Proposition 5.2.
The proof has three steps. Step 1 establishes (5.9). Steps 2 and 3 establish (5.10).
Step 1. We prove (5.9). The stationary distribution of $S$ given $K$ is $N(0,\sigma^{2}(K))$ where $\sigma^{2}(K)=(1-\rho^{2}(K))^{-1}$.
Taking log of the conditional density gives
$\ln f_{S\mid K}(S\mid K)$ $=[(-2\sigma^{2}(K))^{-1}]S^{2}+[-1/2\ln(2\pi\sigma^{2}(K))],$
whose derivative with respect to $K$ is
$\partial_{K}\ln f_{S\mid K}(S\mid K)=\partial_{K}[(-2\sigma^{2}(K))^{-1}]S^{2}+\partial_{K}[-1/2\ln(2\pi\sigma^{2}(K))].$ (B.6)
Since the log joint density is
$\ln f(S,K)=\ln f_{S\mid K}(S\mid K)+\ln f_{K}(K)$
differentiating both sides with respect to $K$ gives
$\partial_{K}\ln f(S,K)=\partial_{K}\ln f_{S\mid K}(S\mid K)+\partial_{K}\ln f_{K}(K).$ (B.7)
Plugging (B.6) into (B.7) gives
$w_{0}(X)=\underbracket{\partial_{K}[(-2\sigma^{2}(K))^{-1}]}_{\gamma_{1}(K)}S^{2}+\underbracket{\partial_{K}[-1/2\ln(2\pi\sigma^{2}(K))-\ln f_{K}(K)]}_{\gamma_{0}(K)}$ (B.8)
which is a linear function of $S^{2}$ with slope $\gamma_{1}(K)$ and intercept $\gamma_{0}(K)$.
Step 2. For any step $k\geq 1$ and any time $t\geq 0$, we show that
$\mathrm{E}[S^{2}_{t}\mid S_{t+k}=S,K]=\dfrac{1-\rho^{2k}(K)}{1-\rho^{2}(K)}+\rho^{2k}(K)S^{2}.$ (B.9)
The proof is established using an inductive argument. In the base case $k=0$, (B.9) holds trivially. To verify the inductive hypothesis for $k=l$, note that
$\mathrm{E}[S_{t}\mid S_{t+1}=S,K]=\rho(K)S,\quad\text{Var}(S_{t}\mid S_{t+1},K)=\sigma^{2}(K)(1-\rho^{2}(K))=1.$
Squaring the first term and adding variance term gives
$\mathrm{E}[S^{2}_{t}\mid S_{t+1},K]=1+\rho^{2}(K)S^{2},$ (B.10)
which is a special case of (B.9) with $k=1$. Next,
$\mathrm{E}[S^{2}_{t}\mid S_{t+l+1}=S]$ $=^{i}\mathrm{E}[\mathrm{E}[S^{2}_{t}\mid S_{t+l},K]\mid S_{t+l+1},K]$
$=^{ii}\mathrm{E}[1+\rho^{2}(K)S^{2}_{t+l}\mid S_{t+l+1}=S,K]$
$=^{iii}1+\rho^{2}(K)\dfrac{1-\rho^{2l}(K)}{1-\rho^{2}(K)}+\rho^{2l+2}(K)S^{2}$
$=^{iv}\dfrac{1-\rho^{2l}(K)+\rho^{2}(K)-\rho^{2(l+1)}(K)}{1-\rho^{2}(K)}+\rho^{2l+2}(K)S^{2},$
where (i) follows from Law of Iterated Expectations, (ii) from the inductive hypothesis (B.9) with $k=l$, (iii) from the inductive step (B.10) at $t=(t+l)$. Simplifying the algebra in (iv) gives an expression that is a special case of (B.9) with $k=(l+1)$.
Step 3. We prove (5.10). Plugging $S^{2}$ into (5.2) gives
$\sum_{t\geq 0}\beta^{t}\mathrm{E}[S^{2}_{-t}\mid S_{0}=S,K]$
$=\dfrac{\sum_{t\geq 0}\beta^{t}}{1-\rho^{2}(K)}+\left(\dfrac{1-\rho^{2t}(K)}{1-\rho^{2}(K)}+\rho^{2t}(K)S^{2}\right)$
$=\dfrac{\sum_{t\geq 0}\beta^{t}}{1-\rho^{2}(K)}-\dfrac{\sum_{t\geq 0}\beta^{t}\rho^{2t}(K)}{1-\rho^{2}(K)}+\sum_{t\geq 0}\beta^{t}\rho^{2t}(K)S^{2}$
$=\dfrac{\sum_{t\geq 0}\beta^{t}}{1-\rho^{2}(K)}-\dfrac{1}{(1-\beta\rho^{2}(K))(1-\rho^{2}(K))}+\dfrac{1}{1-\beta\rho^{2}(K)}S^{2}$
$=\dfrac{1}{1-\beta\rho^{2}(K)}S^{2}+\dfrac{\beta}{1-\beta\rho^{2}(K)}.$
Multiplying the discounted sum above by $\gamma_{1}(K)$ and adding $\gamma_{0}(K)\sum_{t\geq 0}\beta^{t}$ gives (5.10).
$\blacksquare$
Proof of Proposition 7.1.
The statement of Proposition follows from (7.3)–(7.5). Since $\beta<1$, the operator $\mathrm{I}-\mathrm{A}$ is invertible, the minimizer of (7.5) is unique.
$\blacksquare$
Proof of Proposition 7.2.
The statement of Proposition follows from (7.17)–(7.18). Since $\beta<1$, the operator $\mathrm{I}-\mathrm{A}^{*}$ is invertible, the minimizer of (7.22) is unique.
$\blacksquare$
B.1. Proof of Theorem 7.1
Proof of Theorem 7.1.
The proof builds on an argument similar to Lemmas A1–A6 and Theorem 1 in Chernozhukov et al. (2022) (cf Bradic et al. (2022)). The proof of Theorem 7.1 below differs in two respects. First, it generalizes the proof of Theorem 1 by considering projections onto transformed basis $\psi(x)=(\mathrm{I}-\mathrm{A}_{0})b$. Second, it drops Assumption 3 of Chernozhukov et al. (2022).
Step 0. (Notation) Let $\varepsilon_{n}$ be an upper bound obeying
$\|\widehat{G}-G\|_{\infty}=O_{P}(\varepsilon_{n}),\quad\|\widehat{M}-M\|_{\infty}=O_{P}(\varepsilon_{n}),\quad\sqrt{\ln p^{2}/n}=o(\varepsilon_{n}).$ (B.11)
By Assumption 7.1 we can define $J_{0}$ as a set of indices of a sparse approximation with $|J_{0}|=s_{0}$ with $s_{0}\geq C\varepsilon_{n}^{-2/(2\xi+1)}$ and coefficients $\widetilde{\rho}_{j}$ for $j\in J_{0}$ such that for $\widetilde{V}_{0}(x)=\sum_{j\in J_{0}}\widetilde{\rho}_{j}b_{j}(x)$ satisfies
$\mathrm{E}(V_{0}(X)-\widetilde{V}_{0}(X))^{2}\leq Cs^{-2\xi}_{0}$ (B.12)
Define $\psi(x)=(\mathrm{I}-\mathrm{A}_{0})b(x)$ as the vector of projected basis functions. Note that the projection coefficient
$\rho=\arg\min_{v}\|\zeta_{0}-\psi^{\prime}v\|$ (B.13)
can be equivalently written as
$G_{V}\rho-M_{V}=0,$
with $G_{V}=\mathbb{E}[\psi(X)\psi^{\prime}(X)]$ and $M_{V}=\mathbb{E}[\psi(X)\zeta_{0}(X)]$. Define
$\rho_{\ast}\in\arg\min_{v}\;(\rho-v)\prime G_{V}(\rho-v)+2\varepsilon_{n}\sum_{j\in J_{0}^{c}}|v_{j}|\text{.}$ (B.14)
Let $J$ to be the vector of indices of nonzero elements of $\rho_{\ast}$ and $\left|A\right|$ be the number non zero elements of any finite set $A$, and let $J^{c}$ denote $J^{c}=\{1,2,\dots,p\}\setminus J$.
Step 1. We verify the analogs of Lemmas A1–A2 hold with $G_{V}=\mathrm{E}[\psi(X)\psi^{\prime}(X)]$ instead of $\mathrm{E}[b(X)b^{\prime}(X)]$. The statement of Lemma A1
$\|G_{V}(\rho_{\ast}-\rho)\|_{\infty}\leq\varepsilon_{n}$ (B.15)
follows from the definition of $\rho_{\ast}$ and $\rho$. To verify the statement of Lemma A2, note that
$(\rho-\rho_{\ast})^{\prime}G_{V}(\rho-\rho_{\ast})$ $\leq(\rho-\bar{\rho})^{\prime}G_{V}(\rho-\bar{\rho})$
$=\|\psi^{\prime}(\rho-\bar{\rho})\|^{2}_{2}$
$\leq 2\|\psi^{\prime}\rho-\zeta_{0}\|^{2}_{2}+2\|\psi^{\prime}\bar{\rho}-\zeta_{0}\|^{2}_{2}\leq 4\|\psi^{\prime}\bar{\rho}-\zeta_{0}\|^{2}_{2},$
where the last inequality follows from the definition of the projection coefficient $\rho$ in (B.13). As shown in the proof of Theorem E.1, the norm bound
$1-\beta\leq\|\mathrm{I}-\mathrm{A}_{0}\|_{2}\leq 1+\beta.$ (B.16)
implies
$(1+\beta)^{2}\lambda_{\max}(\mathrm{E}[b(X)b(X)^{\prime}])$ $\geq\lambda_{\max}(G_{V})$
$\geq\lambda_{\min}(G_{V})\geq(1-\beta)^{2}\lambda_{\min}(\mathrm{E}[b(X)b(X)^{\prime}])>0$
Plugging $\zeta_{0}=(\mathrm{I}-\mathrm{A}_{0})V_{0}$ and $\psi=(\mathrm{I}-\mathrm{A}_{0})b$ gives
$\|\psi^{\prime}\bar{\rho}-\zeta_{0}\|=\|(\mathrm{I}-\mathrm{A}_{0})(V_{0}-b^{\prime}\bar{\rho})\|\leq\|\mathrm{I}-\mathrm{A}_{0}\|_{2}\|V_{0}-b^{\prime}\bar{\rho}\|_{2}\leq(1+\beta)s_{0}^{-\xi}.$
Combining the bounds and noting that $s_{0}\geq C\varepsilon^{-2/(2\xi+1)}_{n}$
$(\rho-\rho_{\ast})^{\prime}G_{V}(\rho-\rho_{\ast})$ $\leq(\rho-\bar{\rho})^{\prime}G_{V}(\rho-\bar{\rho})\leq 4(1+\beta)^{2}C^{2}\varepsilon^{4\xi/(2\xi+1)}_{n}.$ (B.17)
Step 2. We show that $\|V_{0}-b^{\prime}\rho_{\ast}\|^{2}=O_{P}(\varepsilon^{4\xi/(2\xi+1)}_{n})$. Noting that
$\|b^{\prime}(\bar{\rho}-\rho_{\ast})\|^{2}=\|(\mathrm{I}-\mathrm{A}_{0})^{-1}\psi^{\prime}(\bar{\rho}-\rho_{\ast})\|^{2}$
$\leq^{i}(1-\beta)^{-2}(\bar{\rho}-\rho_{\ast})^{\prime}G(\bar{\rho}-\rho_{\ast})$
$\leq^{ii}2(1-\beta)^{-2}(\bar{\rho}-\rho)^{\prime}G(\bar{\rho}-\rho)+2(1-\beta)^{-2}(\rho-\rho_{\ast})^{\prime}G(\rho-\rho_{\ast})$
$\leq^{iii}8(1-\beta)^{-2}(1+\beta)^{2}C^{2}\varepsilon^{4\xi/(2\xi+1)}_{n}.$
where (i) follows from (B.16) and (ii) from triangle inequality and (iii) from (B.17). Invoking triangle inequality gives
$\|V_{0}-b^{\prime}\rho_{\ast}\|\leq\|V_{0}-b^{\prime}\bar{\rho}\|+\|b^{\prime}(\bar{\rho}-\rho_{\ast})\|=O(\varepsilon^{2\xi/(2\xi+1)}_{n})+O(\varepsilon^{2\xi/(2\xi+1)}_{n})=O(\varepsilon^{2\xi/(2\xi+1)}_{n}),$ (B.18)
where $\|V_{0}-b^{\prime}\bar{\rho}\|=O(\varepsilon^{2\xi/(2\xi+1)}_{n})$ follows from (B.12).
Step 3. We verify Lemmas A3–A4. The statement of Lemma A3 holds with
$|J|\leq C\varepsilon^{-2/(2\xi+1)}$ (B.19)
which follows from Steps 1–3 and $\lambda_{\max}(G_{V})\leq(1+\beta)\bar{\lambda}$. The analogs of Lemmas A4 and A5 also hold.
Step 4. We show that, for any $\Delta\in\mathrm{R}^{p}:\|\Delta_{J^{c}}\|_{1}\leq 3\|\Delta_{J}\|_{1}$, we have
$\Delta^{\prime}G\Delta\leq 2\Delta^{\prime}\widehat{G}\Delta$ (B.20)
For any $\Delta\in\mathrm{R}^{p}$, the properties of the norm imply
$|\Delta^{\prime}G\Delta-\Delta^{\prime}\widehat{G}\Delta|$ $\leq\|\widehat{G}-G\|_{\infty}\|\Delta\|^{2}_{1}.$ (B.21)
Since $\Delta\in\mathrm{R}^{p}$ belongs to restricted set $\|\Delta_{J^{c}}\|_{1}\leq 3\|\Delta_{J}\|_{1}$, additionally,
$\|\Delta\|^{2}_{1}\leq 16\|\Delta_{J}\|^{2}_{1}.$
Next, the term $\|\Delta_{J}\|^{2}_{1}$ can be upper bounded as
$\|\Delta_{J}\|^{2}_{1}$ $\leq|J|\|\Delta_{J}\|^{2}_{2}\leq|J|\lambda^{-1}_{\min}(G)\Delta^{\prime}G\Delta.$
where (i) follows from Cauchy inequality ad (ii) from the definition of $\lambda_{\min}(G)$. Invoking Lemma A3 in Chernozhukov et al. (2022) gives $|J|\leq\varepsilon_{n}^{-2/(2\xi+1)}$ which gives
$|J|\lambda^{-1}_{\min}(G)\Delta^{\prime}G\Delta\leq\varepsilon_{n}^{-2/(2\xi+1)}\lambda^{-1}_{\min}(G)\Delta^{\prime}G\Delta.$
Invoking (B.21) gives a multiplicative error bound
$\Delta^{\prime}G\Delta$ $\leq(1+32\varepsilon_{n}^{(2\xi-1)/(2\xi+1)})\Delta^{\prime}\widehat{G}\Delta\leq 2\Delta^{\prime}\widehat{G}\Delta,$
that matches (B.20).
Step 5. Let $\Delta=\widehat{\rho}-\rho_{\ast}$ be the estimation error. The following upper bound applies
$\Delta^{\prime}G\Delta$ $\leq^{i}2\Delta^{\prime}\widehat{G}\Delta$
$\leq^{ii}6r\|\Delta\|_{1}\leq^{iii}24r\|\Delta_{J}\|_{1}\leq^{iv}24r|J|^{1/2}\|\Delta_{J}\|_{2}$
$\leq^{v}\dfrac{24r}{\lambda^{1/2}_{\min}(G)}|J|^{1/2}(\Delta^{\prime}G\Delta)^{1/2},$
where (i) follows from (B.20), the statements (ii)–(iii) follow the statement in Lemma A5, (iv) follows from Cauchy inequality, and (v) noting that
$\Delta^{\prime}G\Delta\geq\lambda_{\min}(G)\|\Delta\|^{2}\geq\lambda_{\min}(G)\|\Delta_{J}\|^{2}.$
Dividing by $(\Delta^{\prime}G\Delta)^{1/2}$ gives
$\Delta^{\prime}G\Delta=^{i}O_{P}(|J|^{1/2}r)=O_{P}(\varepsilon_{n}^{-1/(2\xi_{V}+1)}\varepsilon_{n}n^{c})=^{ii}O_{P}(\varepsilon_{n}^{(2\xi_{V})/(2\xi_{V}+1)}n^{c}),$
where (i) follows from Step 3 and (ii) from the choice of $r_{V}$ in A5 of Assumption 7.1.
Step 6. We verify that the bound (B.11) holds with $\varepsilon_{n}=n^{-d_{V}}$ where $\sqrt{\ln p_{V}^{2}/n}+a_{n}+\zeta_{n}=o(n^{-d_{V}})$. Define
$\check{G}_{V}$ $=\mathrm{E}_{n}(b(X_{i})-\mathrm{A}_{0}b)(X_{i}))(b(X_{i})-\mathrm{A}_{0}b(X_{i}))^{\prime}$
$\check{M}_{V}$ $=\mathrm{E}_{n}(b(X_{i})-\beta b(X_{i+}))\widehat{\zeta}(X_{i}).$
where $\mathrm{E}_{n}[\cdot]$ is the sample average calculated on respective partition. Decomposing error term gives
$\widehat{G}_{V}-\check{G}_{V}$ $=\mathrm{E}_{n}(b(X_{i})-\mathrm{A}_{0}b(X_{i}))((\mathrm{A}_{0}b-\widehat{A}b)(X_{i}))^{\prime}$
$\mathrm{E}_{n}(b(X_{i})-\mathrm{A}_{0}b(X_{i}))((\mathrm{A}_{0}b-\widehat{A}b)(X_{i}))^{\prime}$
$+\mathrm{E}_{n}((\mathrm{A}_{0}b-\widehat{A}b)(X_{i}))((\mathrm{A}_{0}b-\widehat{A}b)(X_{i}))^{\prime}$
$\widehat{M}_{V}-\check{M}_{V}$ $=\mathrm{E}_{n}(b(X_{i})-\beta b(X_{i+}))(\zeta_{0}(X_{i})-\widehat{\zeta}(X_{i})).$
By Law of Large Numbers, the following concentration bounds hold for sample averages
$\|\check{G}_{V}-G_{V}\|_{\infty}=o_{P}(\sqrt{\ln p_{V}^{2}/n}),\quad\|\check{M}_{V}-M_{V}\|_{\infty}=o_{P}(\sqrt{\ln p_{V}^{2}/n}).$
as well as for demeaned error terms
$\|\widehat{G}_{V}-\check{G}_{V}\|_{\infty}=o_{P}(n^{-d_{V}}),\quad\|\widehat{M}_{V}-\check{M}_{V}\|_{\infty}=o_{P}(n^{-d_{V}}).$
Step 7. (Conclusion) The upper bound can be seen as
$\|\widehat{V}-V_{0}\|^{2}$ $\leq 2\|V_{0}-b^{\prime}\rho_{\ast}\|^{2}+2\|b^{\prime}(\widehat{\rho}-\rho_{\ast})\|^{2}$
$\leq 2Cs_{0}^{-2\xi}+\Delta^{\prime}(\mathrm{E}[b(X)b(X)^{\prime}])\Delta.$
The approximation error term is bounded by $s_{V}^{-2\xi_{V}}\leq Cn^{-2d_{V}\xi_{V}/(2\xi_{V}+1)}$ as assumed in Assumption 7.1. The second term is bounded as
$\Delta^{\prime}(\mathrm{E}[b(X)b(X)^{\prime}])\Delta$
$\leq\bar{\lambda}\Delta^{\prime}\Delta\leq(\bar{\lambda}/\underline{\lambda})\Delta^{\prime}G\Delta=O_{P}(\varepsilon_{n}^{2(2\xi)/(2\xi+1)}n^{c}).$
$\blacksquare$
Proof of Theorem 7.2.
We focus on the case when $m(\cdot,\cdot)$ is known. By Law of Large Numbers, the following concentration bounds hold for sample averages
$\|\widehat{G}^{\alpha}-G^{\alpha}\|_{\infty}=o_{P}(\sqrt{\ln p_{\alpha}^{2}/n}),\quad\|\widehat{M}^{\alpha}-M^{\alpha}\|_{\infty}=o_{P}(n^{-d_{\alpha}}).$
which implies $\varepsilon_{n}=n^{-d_{\alpha}}$ where $\sqrt{\ln p^{2}/n}+a^{*}_{n}=o(n^{-d_{\alpha}})$ suffices for
$\|\widehat{G}^{\alpha}-G^{\alpha}\|_{\infty}=O_{P}(\varepsilon_{n}),\quad\|\widehat{M}^{\alpha}-M^{\alpha}\|_{\infty}=O_{P}(\varepsilon_{n}).$ (B.22)
The rest of the proof follows similarly to the proof of Theorem 7.1.
$\blacksquare$
Appendix C Asymptotic Theory
C.1. Time-Invariant Case.
In this section we derive sufficient conditions for the estimator $\widehat{\delta}$ to be asymptotically linear.
$\sqrt{n}(\widehat{\delta}-\delta_{0})$ $=\frac{1}{n}\sum_{\ell=1}^{L}\sum_{i\in I_{\ell}}[(1-\beta)^{-1}\widehat{w}_{\ell}(K_{i})\widehat{\zeta}_{\ell}(X_{i})+\widehat{\phi}_{w\ell}(Z_{i})+\widehat{\phi}_{\zeta\ell}(Z_{i})+o_{P}(1).$ (C.1)
Assumption C.1 is a mild consistency condition.
Assumption C.1.
(1)For any partition index $\ell=1,2,\dots,L$ the following convergence conditions hold:
  1. A
    $\int(\widehat{\phi}_{\zeta\ell}(z)-\phi_{\zeta}(z))^{2}F_{0}(dz)=o_{P}(1)$
  2. B
    $\int(\widehat{\phi}_{m\ell}(z)-\phi_{m}(z))^{2}F_{0}(dz)=o_{P}(1)$
  3. C
    $\|\widehat{\zeta}_{\ell}-\zeta_{0}\|=o_{P}(1)$.
(2) For any partition index $\ell=1,2,\dots,L$, $\|\widehat{w}_{\ell}-w\|=o_{P}(1)$.
Assumption C.2 is a small bias condition that controls higher order bias of plug-in estimators.
Assumption C.2.
For any partition index $\ell=1,2,\dots,L$ the following higher-order bias terms are negligible
$n^{1/2}\int_{\mathcal{Z}}[(1-\beta)^{-1}w_{0}(k)(\widehat{\zeta}_{\ell}(x)-\zeta_{0}(x))+\widehat{\phi}_{\zeta\ell}(z)-\phi_{\zeta}(z)]F_{0}(dz)=o_{P}(1)$
$n^{1/2}\int_{\mathcal{Z}}[(1-\beta)^{-1}(\widehat{w}_{\ell}(k)-w_{0}(k))\zeta_{0}(x)+\widehat{\phi}_{w\ell}(z)-\phi_{w}(z)]F_{0}(dz)=o_{P}(1)$
Assumption C.3 includes a product rate condition as well as the technical conditions on the estimators of $w$ and $\zeta$.
Assumption C.3.
For any partition index $\ell=1,2,\dots,L$ (1) the following products of first-stage errors converge fast enough
$n^{1/2}\|\widehat{w}_{\ell}-w_{0}\|\|\widehat{\zeta}_{\ell}-\zeta_{0}\|=o_{P}(1).$
(2) The estimate $\widehat{w}_{\ell}(k)$, as well as the weight function $w_{0}(k)$, and the estimate $\widehat{\zeta}_{\ell}(x)$, as well as the function $\zeta_{0}(x)$, are uniformly bounded over the support of $\mathcal{X}$.
Lemma C.1 (Asymptotic Theory for Time-Invariant Case).
If Assumptions C.1C.3 hold then (C.1) holds, and the asymptotic variance estimate is consistent
$\widehat{\Omega}\rightarrow^{p}\Omega.$ (C.2)
Proof of Lemma C.1.
Throughout this proof, let $C>0$ denote a generic constant (possibly different each time it appears).
Step 1 . Let $\ell\in\{1,2,\dots,L\}$ indicate the partition index. Define the following error terms
$R_{1,\ell}(Z)$ $=(1-\beta)^{-1}w_{0}(K)(\widehat{\zeta}_{\ell}(X)-\zeta_{0}(X))+\widehat{\phi}_{\zeta\ell}(Z)-\phi_{\zeta}(Z)$
$R_{2,\ell}(Z)$ $=(1-\beta)^{-1}(\widehat{w}_{\ell}(K)-w_{0}(K))\zeta_{0}(X)+\widehat{\phi}_{w\ell}(Z)-\phi_{w}(Z)$
$R_{3,\ell}(Z)$ $=(1-\beta)^{-1}(\widehat{w}_{\ell}(K)-w_{0}(K))(\widehat{\zeta}_{\ell}(X)-\zeta_{0}(X))$
and decompose the error term as
$\psi(Z_{i},\widehat{\gamma}_{\ell},\widehat{\phi}_{\ell})-\psi(Z_{i},\gamma_{0},\phi_{0})=R_{1,\ell}(Z_{i})+R_{2,\ell}(Z_{i})+R_{3,\ell}(Z_{i}).$
Step 2 shows
$n^{-1/2}[\sum_{i=1}^{n}R_{2,\ell}(Z_{i})+R_{3,\ell}(Z_{i})]=o_{P}(1).$ (C.3)
Step 3 shows
$n^{-1/2}\sum_{i=1}^{n}R_{1,\ell}(Z_{i})=o_{P}(1).$ (C.4)
Step 4 shows (C.2).
Step 2 (Proof of (C.3)). By Assumption C.1, we have $n^{1/2}\mathrm{E}[R_{2,\ell}(Z_{i})+R_{1,\ell}(Z_{i})]=o_{P}(1)$. Furthermore, we have
$\mathrm{E}[R^{2}_{2,\ell}(Z)]\leq 2\mathrm{E}(\widehat{\phi}_{w\ell}(Z)-\phi_{w}(Z))^{2}+2(1-\beta)^{-2}C\mathrm{E}(\widehat{w}_{\ell}(K)-w_{0}(K))^{2}=o_{P}(1)$
and
$\mathrm{E}[R^{2}_{1,\ell}(Z)]\leq 2\mathrm{E}(\widehat{\phi}_{\zeta\ell}(Z)-\phi_{\zeta}(Z))^{2}+2(1-\beta)^{-2}C\mathrm{E}(\widehat{\zeta}_{\ell}(X)-\zeta_{0}(X))^{2}=o_{P}(1)$
where $C$ is a uniform bound in Assumption C.3. Collecting the terms gives (C.1).
Step 3 (Proof of (C.4)) . We show (C.4). By Assumption C.3 (1),
$n^{1/2}\mathrm{E}[R_{1,\ell}(Z)]$ $=n^{1/2}\int_{\mathcal{Z}}(\widehat{w}_{\ell}(k)-w_{0}(k))(\widehat{\zeta}_{\ell}(x)-\zeta_{0}(x))F_{0}(dx)$
$\leq n^{1/2}\|\widehat{w}_{\ell}-w_{0}\|\cdot\|\widehat{\zeta}_{\ell}-\zeta_{0}\|=o_{P}(1).$
By Assumptions C.1 (3)-(4) and C.3 (2), the term $\mathrm{E}[R^{2}_{1,\ell}(Z)]$ is bounded as
$\mathrm{E}[R^{2}_{1,\ell}(Z)]\leq 2C(1-\beta)^{-1}\min(\|\widehat{w}_{\ell}-w_{0}\|,\|\widehat{\zeta}_{\ell}-\zeta_{0}\|)=o_{P}(1)$
where $C$ is a uniform bound in Assumption C.3. Steps 1–3 imply that (C.1) holds, and, therefore, $\widehat{\delta}\rightarrow^{p}\delta_{0}$.
Step 4 (Proof of (C.2)). Assumption C.1 gives
$\int_{\mathcal{Z}}(\widehat{w}_{\ell}(k)\widehat{\zeta}_{\ell}(x)-w_{0}(k)\zeta_{0}(x))^{2}F_{0}(dz)$
$\leq 2\int_{\mathcal{Z}}(\widehat{w}_{\ell}(k)\widehat{\zeta}_{\ell}(x)-w_{0}(k)\widehat{\zeta}_{\ell}(x))^{2}F_{0}(dz)+\int_{\mathcal{Z}}(w_{0}(k)(\widehat{\zeta}_{\ell}(x)-\zeta_{0}(x)))^{2}F_{0}(dz)$
$\leq C\|\widehat{w}_{\ell}-w_{0}\|^{2}+C\|\widehat{\zeta}_{\ell}-\zeta_{0}\|^{2}.$
Decomposing the error of the original moment gives
$\int_{\mathcal{Z}}((1-\beta)^{-1}\widehat{w}_{\ell}(k)\widehat{\zeta}_{\ell}(x)-(1-\beta)^{-1}w_{0}(k)\zeta_{0}(x)-\widehat{\delta}+\delta_{0})^{2}F_{0}(dz)$
$\leq C\|\widehat{w}_{\ell}-w_{0}\|^{2}+C\|\widehat{\zeta}_{\ell}-\zeta_{0}\|^{2}+C\int_{\mathcal{Z}}(\widehat{\delta}-\delta_{0})^{2}F_{0}(dz)=o_{P}(1)+o_{P}(1)=o_{P}(1).$
Invoking Lemma 16 in CEINR gives $\widehat{\Omega}\rightarrow^{p}\Omega$. $\blacksquare$
In this Section we derive sufficient conditions for the estimator $\widehat{\delta}$ of $\delta_{0}$ to be asymptotically linear, that is
$\sqrt{n}(\widehat{\delta}-\delta_{0})$ $=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}[m_{0}(Z_{i},V_{0})-\delta_{0}+\alpha_{0}(X_{i})\lambda(Z_{i},V_{0})+\phi_{\zeta}(Z_{i})+\phi_{w}(Z_{i})]+o_{P}(1).$ (C.5)
Assumption C.4.
For any partition index $\ell=1,2,\dots,L$ the following higher-order bias terms are negligible
$n^{1/2}\int_{\mathcal{Z}}[\alpha_{0}(x)(\widehat{\zeta}_{\ell}(x)-\zeta_{0}(x))+\widehat{\phi}_{\zeta\ell}(z)-\phi_{\zeta}(z)]F_{0}(dz)=o_{P}(1)$
$n^{1/2}\int_{\mathcal{Z}}[\widehat{m}_{\ell}(z,V_{0})-m_{0}(z,V_{0})+\widehat{\phi}_{m\ell}(z)-\phi_{m}(z)]F_{0}(dz)=o_{P}(1)$
Assumption C.5.
For any partition index $\ell=1,2,\dots,L$ (1) the following products of first-stage errors converge fast enough
$n^{1/2}(\|\widehat{V}_{\ell}-V_{0}\|+\|\widehat{\zeta}_{\ell}-\zeta_{0}\|)\|\widehat{\alpha}_{\ell}-\alpha_{0}\|=o_{P}(1).$
(2) the function $\alpha_{0}(x)$ and its estimate $\widehat{\alpha}_{\ell}(x)$ are uniformly bounded over the support of $\mathcal{X}$ by some constant $C_{\alpha}$.
(3) the function $\mathrm{E}[m_{0}^{2}(Z,V)]\leq C_{m}\|V\|^{2}$ for some finite $C_{m}$ and for any integrable function $V$.
Assumption C.6.
For any partition index $\ell=1,2,\dots,L$ (1) the following condition holds
$n^{1/2}\int_{\mathcal{Z}}(\widehat{m}_{\ell}(z,\widehat{V}_{\ell})-m_{0}(z,\widehat{V}_{\ell})-\widehat{m}_{\ell}(z,V_{0})+m_{0}(z,V_{0}))F_{0}(dz)=o_{P}(1).$ (C.6)
Example C.1 (continues=ex:avder).
Since $m(z,V)=\partial_{K}V(X)$ is a known functional of $V$, Assumption C.6 is automatically satisfied.
Lemma C.2 (Asymptotic Theory for Time-Variant Case).
If Assumptions C.1(1), C.3, C.4, C.5, C.6 hold, then (C.5) holds, and the asymptotic variance estimate is consistent $\widehat{\Omega}\rightarrow^{p}\Omega$.
Proof of Lemma C.2.
We verify Assumptions 1–3 of (Chernozhukov et al. (2022), CEINR) where notation is redefined as follows. When writing the residual $\lambda(Z,V)$, we make the dependence on $\zeta$ in $\lambda(Z,V)$ explicit, that is,
$\lambda(Z,V,\zeta)=\beta V(X_{+})-V(X)+\zeta(X).$
The following notation mapping from CEINR to this paper is used
$\theta_{CNR}$ $:=\delta,$
$\gamma_{CNR}$ $:=(m(\cdot,\cdot),V(\cdot),\zeta(\cdot)),$
$w_{CNR}$ $=z$
$g_{CNR}(w_{CNR},\gamma_{CNR},\theta_{CNR})$ $=m(z,V)-\delta$
$\alpha_{CNR}$ $=(\alpha,\phi_{\zeta},\phi_{m})$
Note that the correction term does not depend on the target parameter, that is,
$\phi_{CNR}(w_{CNR},\gamma_{CNR},\alpha_{CNR},\theta_{CNR})$ $=\phi_{CNR}(w_{CNR},\gamma_{CNR},\alpha_{CNR})$
$=\alpha(x)\lambda(z,V,\zeta)+\phi_{\zeta}(z)+\phi_{m}(z)$
Steps 1, 2, 3 verify Assumptions 1,2,3, respectively. Step 4 establishes (C.2).
Step 1 . The term in Assumption 1(ii) of CEINR is bounded as
$\int_{\mathcal{W}_{CNR}}(\phi_{CNR}(w_{CNR},\alpha_{0,CNR},\widehat{\gamma}_{\ell,CNR})-\phi_{CNR}(w_{CNR},\alpha_{0,CNR},\gamma_{0,CNR}))^{2}F_{0}(dw_{CNR})$
$\leq\int_{\mathcal{Z}}\alpha^{2}_{0}(x)\lambda^{2}(z,\widehat{V}_{\ell}-V_{0},\widehat{\zeta}_{\ell}-\zeta_{0})F_{0}(dz)$
[…]
The term in Assumption 1(iii) of CEINR is bounded as
$\int_{\mathcal{W}_{CNR}}(\phi_{CNR}(w_{CNR},\widehat{\alpha}_{\ell,CNR},\gamma_{0,CNR})-\phi_{CNR}(w_{CNR},\alpha_{0,CNR},\gamma_{0,CNR}))^{2}F_{0}(dw_{CNR})$
$\leq 3\underbracket{\int_{\mathcal{Z}}(\widehat{\phi}_{m\ell}(z)-\phi_{m}(z))^{2}F_{0}(dz)}_{T_{1\ell}}+3\underbracket{\int_{\mathcal{Z}}(\widehat{\phi}_{\zeta\ell}(z)-\phi_{\zeta}(z))^{2}F_{0}(dz)}_{T_{2\ell}}$
$+3\underbracket{\int_{\mathcal{Z}}(\widehat{\alpha}_{\ell}(x)-\alpha_{0}(x))^{2}\lambda^{2}(z,V_{0},\zeta_{0})F_{0}(dz)}_{T_{3\ell}}=3\sum_{j=1}^{3}T_{j\ell}.$
Invoking Assumption C.1 ensures that $T_{1\ell}+T_{2\ell}=o_{P}(1)$. To bound $T_{3\ell}$, we have
$T_{3\ell}\leq[\sup_{x\in\mathcal{X}}\mathrm{E}[\lambda^{2}(Z,V_{0},\zeta_{0})\mid X=x]]\int_{\mathcal{Z}}(\widehat{\alpha}_{\ell}(x)-\alpha_{0}(x))^{2}F_{0}(dx)$
which implies $T_{3\ell}=o_{P}(1)$.
Finally, we verify Assumption 1(i) of CEINR. By Assumption C.5(3),
$\int_{\mathcal{Z}}(m_{0}(z,\widehat{V}_{\ell})-m_{0}(z,V_{0}))^{2}F_{0}(dz)\leq C_{m}\int_{\mathcal{Z}}\|\widehat{V}_{\ell}-V_{0}\|^{2}F_{0}(dz)$
Likewise, by Assumption C.6 with $\Delta_{m}(\cdot)$ defined therein
$\int_{\mathcal{Z}}(\widehat{m}_{\ell}(z,\widehat{V}_{\ell})-m_{0}(z,\widehat{V}_{\ell}))^{2}F_{0}(dz)$
$\leq 2\int_{\mathcal{Z}}(\widehat{m}_{\ell}(z,V_{0})-m_{0}(z,V_{0}))^{2}F_{0}(dz)+2\int_{\mathcal{Z}}\Delta^{2}_{m\ell}(z)F_{0}(dz)$
$\leq 2C_{m}\int_{\mathcal{Z}}\|\widehat{V}_{\ell}-V_{0}\|^{2}F_{0}(dz)+2\int_{\mathcal{Z}}\Delta^{2}_{m\ell}(z)F_{0}(dz)=o_{P}(1)+o_{P}(1).$
Combining two bounds gives an upper bound
$\int_{\mathcal{W}_{CNR}}(g_{CNR}(w_{CNR},\widehat{\gamma}_{\ell,CNR},\theta_{0,CNR})-g_{CNR}(w_{CNR},\gamma_{0,CNR},\theta_{0,CNR}))^{2}F_{0}(dw_{CNR})$
$=\int_{\mathcal{Z}}(\widehat{m}_{\ell}(z,\widehat{V}_{\ell})-m_{0}(z,V_{0}))^{2}F_{0}(dz)$
$\leq 2\int_{\mathcal{Z}}(\widehat{m}_{\ell}(z,\widehat{V}_{\ell})-m_{0}(z,\widehat{V}_{\ell}))^{2}F_{0}(dz)+2\int_{\mathcal{Z}}(m_{0}(z,\widehat{V}_{\ell})-m_{0}(z,V_{0}))^{2}F_{0}(dz)=o_{P}(1).$
Step 2 . We verify Assumption 2(i) of CEINR. Let $\ell\in\{1,2,\dots,L\}$ indicate the partition index and $\Delta_{V}(X)=V(X)-V_{0}(X)$. Define the following error terms
$L_{1,\ell}(Z)$ $=\beta(\widehat{\alpha}_{\ell}(X)-\alpha_{0}(X))\Delta_{\widehat{V}_{\ell}}(X_{+})$
$L_{2,\ell}(Z)$ $=-(\widehat{\alpha}_{\ell}(X)-\alpha_{0}(X))\Delta_{\widehat{V}_{\ell}}(X)$
$L_{3,\ell}(Z)$ $=(\widehat{\alpha}_{\ell}(X)-\alpha_{0}(X))(\widehat{\zeta}_{\ell}(X)-\zeta_{0}(X)).$
Decomposing the error term in Assumption 2(i) gives
$\widehat{\Delta}_{CNR,\ell}(W_{CNR})=\sum_{j=1}^{3}L_{j,\ell}(Z).$
Note that
$n^{1/2}|\mathrm{E}[L_{1,\ell}(Z)]|+n^{1/2}\mathrm{E}[L_{2,\ell}(Z)]|$ $\leq 2n^{1/2}\|\widehat{\alpha}_{\ell}-\alpha_{0}\|\|\widehat{V}_{\ell}-V_{0}\|=o_{P}(1)$
$n^{1/2}|\mathrm{E}[L_{3,\ell}(Z)]|$ $\leq n^{1/2}\|\widehat{\alpha}_{\ell}-\alpha_{0}\|\|\widehat{\zeta}_{\ell}-\zeta_{0}\|.$
Finally, since $\sup_{x\in\mathcal{X}}\widehat{\alpha}_{\ell}(x)$ is bounded a.s., we have
$\sup_{j\in\{1,2\}}\mathrm{E}[L^{2}_{j,\ell}(Z)]\leq 2C_{\alpha}^{2}\|\widehat{V}_{\ell}-V_{0}\|^{2}=o_{P}(1).$
For the third term,
$\mathrm{E}[L^{2}_{3,\ell}(Z)]\leq 2C^{2}\|\widehat{\zeta}_{\ell}-\zeta_{0}\|^{2}=o_{P}(1).$
Step 3 . We verify Assumption 3(iv) of CEINR. Let $\ell\in\{1,2,\dots,L\}$ indicate the partition index. Define the following error terms
$B_{1,\ell}$ $=n^{1/2}\mathrm{E}[\widehat{m}_{\ell}(Z,V_{0})-m(Z,V_{0})+\widehat{\phi}_{m\ell}(Z)-\phi_{m}(Z)]$
$B_{2,\ell}:$ $=n^{1/2}\mathrm{E}[\alpha_{0}(X)(\widehat{\zeta}_{\ell}(X)-\zeta_{0}(X))+\widehat{\phi}_{\zeta\ell}(Z)-\phi_{\zeta}(Z)]$
$B_{3,\ell}:$ $=n^{1/2}\mathrm{E}[\widehat{m}_{\ell}(Z,\widehat{V}_{\ell})-m_{0}(Z,\widehat{V}_{\ell})-\widehat{m}_{\ell}(Z,V_{0})+m_{0}(Z,V_{0})]$
$B_{4,\ell}:$ $=\mathrm{E}[w_{0}(X)\Delta_{\widehat{V}_{\ell}}(X)+\alpha_{0}(X)\lambda(Z,\widehat{V}_{\ell}-V_{0})].$
Notice that
$n^{1/2}\mathrm{E}[\widehat{m}(Z,\widehat{V}_{\ell})-m(Z,V_{0})+\alpha_{0}(X)\lambda(Z,\widehat{V}_{\ell}-V_{0},\widehat{\zeta}_{\ell}-\zeta_{0})$
$+\widehat{\phi}_{\zeta}(Z)-\phi_{\zeta}(Z)+\widehat{\phi}_{m\ell}(Z)-\phi_{m}(Z)]=\sum_{j=1}^{4}B_{j,\ell}.$
By Assumption C.4, $B_{1,\ell}+B_{2,\ell}=o_{P}(1)$; By Assumption C.6, $B_{3,\ell}=o_{P}(1)$. Lemma 5.1 implies $B_{4,\ell}=0$, which verifies Assumption 3(iv). Assumption 3 (i) is automatically satisfied since
$\int_{\mathcal{Z}}\phi_{CNR}(w_{CNR},\widehat{\alpha}_{\ell,CNR},\gamma_{0,CNR})F_{0}(dw_{CNR})$
$=\int_{\mathcal{Z}}\widehat{\alpha}_{\ell}(x)\lambda(z,V_{0})F_{0}(dz)=0.$
Likewise, by definition of $\phi_{\zeta}$ and $\phi_{m}$ as correction terms, $\int_{\mathcal{Z}}\widehat{\phi}_{\ell m}(z)F_{0}(dz)+\int_{\mathcal{Z}}\widehat{\phi}_{\ell\zeta}(z)F_{0}(dz)=0$.
Step 4 . Finally, by the first conclusion, the estimator is asymptotically linear, and therefore consistent $\widehat{\delta}\rightarrow^{p}\delta_{0}$. We verifying the remaining conditions of Lemma 16 in CEINR. Decomposing the error of non-orthogonal moment gives
$\int_{\mathcal{Z}}(\widehat{m}_{\ell}(z,\widehat{V}_{\ell})-m_{0}(z,V_{0})-\widehat{\delta}+\delta_{0})^{2}F_{0}(dz)$
$\leq 2\int_{\mathcal{Z}}(\widehat{m}_{\ell}(z,\widehat{V}_{\ell})-m_{0}(z,V_{0}))^{2}F_{0}(dz)+2\int_{\mathcal{Z}}(\widehat{\delta}-\delta_{0})^{2}F_{0}(dz)=o_{P}(1)+o_{P}(1)=o_{P}(1).$
The final condition has been verified in Step 2, which implies $\widehat{\Omega}\rightarrow^{p}\Omega$.
$\blacksquare$
Appendix D First-Stage Rates for Estimators
D.1. General Case Extremum Estimators
In this Section, we give a general statement for plug-in extremum estimators with a nuisance component, nesting Theorems E.1 and E.2 as special cases. Abusing notation, let $\ell(z,V,\omega):\mathcal{Z}\times\mathcal{V}\times\Omega\rightarrow\mathrm{R}$ denote an $M$-estimator loss function whose arguments are the data vector $z$, the scalar function $V$, and the vector-valued nuisance parameter $\omega:\mathcal{Z}\rightarrow\mathrm{R}^{d_{\omega}}$ whose true value $\omega_{0}$ is identified. Define the true value $V_{0}$ as
$V_{0}$ $=\arg\min L(V,\omega_{0}).$ (D.1)
and
$V^{\ast}=\arg\min_{V\in\mathcal{V}_{n}}L(V,\omega_{0}).$
and a plug-in estimate
$\widehat{V}=\arg\min_{\mathcal{V}_{n}}L^{V}_{n}(V)$ (D.2)
where the parameter space $\mathcal{V}$ can be approximated by a growing sieve space $\{\mathcal{V}_{n}\}_{n\geq 1}$ and $\widehat{\omega}$ is estimated on auxiliary sample and $L^{V}_{n}(V)=n^{-1}\sum_{i=1}^{n}\ell(Z_{i},V,\widehat{\omega})$. The results in this section build on earlier work by Nekipelov et al. (2022), Foster and Syrgkanis (2019), and Chernozhukov et al. (2024).
Assumption D.1.
For some sequence $\omega_{n}=o(1)$, the rate bound holds $\|\widehat{\omega}-\omega_{0}\|_{1}=O_{P}(\omega_{n})$ where $\|\omega\|_{1}=\sum_{j=1}^{d_{\omega}}\|\omega_{j}(Z)\|_{Z\sim P_{Z}}$.
Define
$\mathrm{star}(\mathcal{V}_{n}-V_{0})=~{}$$\{x\rightarrow t\,(V(x)-V_{0}(x)):V\in\mathcal{V}_{n},\text{ }t\in[0,1]\}.$
Let $\delta_{n}$ be an upper bound on the critical radius of $\mathrm{star}(\mathcal{V}_{n}-V_{0})$ and all other mean square continuous transformations of $V$.
Assumption D.2.
(1) (Identification) For some positive $\bar{\lambda}\geq\underline{\lambda}>0$, the following lower bound holds
$L(V,\omega_{0})-L(V_{0},\omega_{0})\geq\underline{\lambda}\|V-V_{0}\|^{2},\quad\text{ for any }V\in\mathcal{V},$ (D.3)
and the upper bound holds
$\bar{\lambda}\|V-V_{0}\|^{2}\geq L(V,\omega_{0})-L(V_{0},\omega_{0}).$ (D.4)
(2) (Lipshitz in $(V,\omega)$) For some positive finite constant $C_{L}$,
$|L(V,\omega)-L(V_{0},\omega)-L(V,\omega_{0})+L(V_{0},\omega_{0})|\leq C_{L}\|V-V_{0}\|\|\omega-\omega_{0}\|_{1}\text{ for any }V,\omega.$ (D.5)
(3) For each $\omega$ in the realization set $\Omega$, the function $\ell(z,V,\omega)$ is $L$-Lipshitz in some mean square continuous (possibly, vector-valued) transformation of $V$ whose critical radius is upper bounded by $\delta_{n}$. Specifically, the following condition holds: with probability $1-\iota$,
$L(V,\omega)-L(V^{\ast},\omega)$ $\leq L_{n}(V,\omega)-L_{n}(V^{\ast},\omega)+\delta\|V-V^{\ast}\|+\delta^{2},$ (D.6)
where $\delta=\delta_{n}+c_{0}\sqrt{\ln(c_{1}/\iota)/n}$ for some $c_{0},c_{1}>0$.
Theorem D.1.
Suppose Assumptions D.1 and D.2 hold. Then, with probability $1-\iota$, for some absolute constant $C$
$\left\|\widehat{V}-V_{0}\right\|^{2}\leq C\left(\delta_{n}^{2}+\omega^{2}_{n}+\|V^{\ast}-V_{0}\|^{2}+\frac{\ln(1/\iota)}{n}\right).$ (D.7)
Proof of Theorem D.1.
Step 1 shows
$\underline{\lambda}\|\widehat{V}-V^{\ast}\|^{2}-C_{L}\|\widehat{\omega}-\omega_{0}\|_{1}\|\widehat{V}-V_{0}\|\leq\delta\|\widehat{V}-V^{\ast}\|+\delta^{2}+2\bar{\lambda}\|V^{\ast}-V_{0}\|^{2}.$ (D.8)
Step 2 shows (D.7).
Step 1. Invoking triangular inequality gives
$\|V-V^{\ast}\|^{2}-\|V_{0}-V^{\ast}\|^{2}\leq\|V-V_{0}\|^{2}$ (D.9)
For any $V\in\mathcal{V}$ and the true value $V_{0}$,
$\underline{\lambda}\|V-V_{0}\|^{2}-C_{L}\|\omega-\omega_{0}\|_{1}\|V-V_{0}\|$
$\leq^{(i)}L(V,\omega_{0})-L(V_{0},\omega_{0})-C_{L}\|\omega-\omega_{0}\|_{1}\|V-V_{0}\|$
$\leq^{(ii)}L(V,\omega)-L(V_{0},\omega)$
where (i) from Assumption D.2(1) and (ii) from Assumption D.2(2). Adding and subtracting approximation term $L(V^{\ast},\omega)$ gives
$L(V,\omega)-L(V_{0},\omega)$ $=L(V,\omega)-L(V^{\ast},\omega)+L(V^{\ast},\omega)-L(V_{0},\omega)$
$\leq^{(iii)}L(V,\omega)-L(V^{\ast},\omega)+\bar{\lambda}\|V^{\ast}-V_{0}\|^{2}$
where (iii) follows from (D.4). Plugging $V:=\widehat{V}$ and $\omega:=\widehat{\omega}$ into (D.6) and invoking Assumption D.2(3) and
$L_{n}(\widehat{V},\widehat{\omega})-L_{n}(V^{\ast},\widehat{\omega})\leq 0,$
gives (D.8).
Step 2. To show (D.7), we replace product terms $\delta\|\widehat{V}-V^{\ast}\|$ and $C_{L}\|\omega-\omega_{0}\|_{1}\|V-V_{0}\|$ by upper bounds. Invoking AM-GM inequality for $\delta\|\widehat{V}-V^{\ast}\|$ gives
$\delta\|\widehat{V}-V^{\ast}\|$ $\leq\underline{\lambda}/4\|\widehat{V}-V^{\ast}\|^{2}+\delta^{2}/\underline{\lambda}.$
Invoking AM-GM inequality
$C_{L}\|\widehat{\omega}-\omega_{0}\|_{1}\|\widehat{V}-V_{0}\|\leq\underline{\lambda}/4\|\widehat{V}-V_{0}\|^{2}+C^{2}_{L}/\underline{\lambda}\|\widehat{\omega}-\omega_{0}\|^{2}_{1}.$
Combining (D.8) with AM-GM upper bounds gives
$3/4\underline{\lambda}\|\widehat{V}-V^{\ast}\|^{2}\leq 2\bar{\lambda}\|V^{\ast}-V_{0}\|^{2}+\underline{\lambda}/4\|\widehat{V}-V_{0}\|^{2}+C^{2}_{L}/\underline{\lambda}\|\widehat{\omega}-\omega_{0}\|^{2}_{1}+\delta^{2}(1+1/\underline{\lambda}).$
Adding $3/4\underline{\lambda}\|V^{\ast}-V_{0}\|^{2}$ to each side and invoking
$3/8\|\widehat{V}-V_{0}\|^{2}\leq 3/4\|\widehat{V}-V^{\ast}\|^{2}+3/4\|V^{\ast}-V_{0}\|^{2}$
gives
$3/8\|\widehat{V}-V_{0}\|^{2}\leq 3\bar{\lambda}\|V^{\ast}-V_{0}\|^{2}+\underline{\lambda}/4\|\widehat{V}-V_{0}\|^{2}+C^{2}_{L}/\underline{\lambda}\|\widehat{\omega}-\omega_{0}\|^{2}_{1}+\delta^{2}(1+1/\underline{\lambda}).$
which implies
$\underline{\lambda}/8\|\widehat{V}-V_{0}\|^{2}\leq 3\bar{\lambda}\|V^{\ast}-V_{0}\|^{2}+C^{2}_{L}/\underline{\lambda}\|\widehat{\omega}-\omega_{0}\|^{2}_{1}+\delta^{2}(1+1/\underline{\lambda}).$
Dividing by $\underline{\lambda}/8$ gives
$\|\widehat{V}-V_{0}\|^{2}\leq 24\bar{\lambda}/\underline{\lambda}\|V^{\ast}-V_{0}\|^{2}+8C^{2}_{L}/\underline{\lambda}^{2}\|\widehat{\omega}-\omega_{0}\|^{2}_{1}+8\delta^{2}(1/\underline{\lambda}+1/\underline{\lambda}^{2}).$
which matches (D.7) with $C\geq\max(24\bar{\lambda}/\underline{\lambda},8C^{2}_{L}/\underline{\lambda}^{2},8/\underline{\lambda},8/\underline{\lambda}^{2})$ on the event $\|\widehat{\omega}-\omega_{0}\|^{2}_{1}\leq\omega^{2}_{n}$.
$\blacksquare$
Appendix E First-Stage Rates for Neural Network Estimators of Value Function and Dynamic Riesz Representer
In this Section, we derive mean square convergence rates for estimators of the value function and dynamic Riesz representer based on Theorem D.1. Let $\ell_{V}(z,V,\omega)$ denote an $M$-estimator loss function as in (7.8).
Assumption E.1.
(A) There exists a vanishing numeric sequence $\zeta_{n}=o(1)$ for which $\|\widehat{\zeta}-\zeta_{0}\|=O_{P}(\zeta_{n})$.
(B) There exists a vanishing numeric sequence $a_{n}=o(1)$ for which $\|\widehat{A}-A_{0}\|=O_{P}(a_{n})$.
Assumption E.1 is a basic consistency condition for the plug-in estimators of per-period utility $\zeta$. For the dynamic binary choice in Section A, the per-period utility $\zeta(x)$ is a known smooth transformation of the conditional choice probability and a structural parameter given in (2.4)–(A.3).
Assumption E.2.
The estimated operator $\widehat{A}$ belongs to a realization set $\mathcal{\mathrm{A}}_{n}$ with probability $1-o(1)$. For some constant $C_{A}$, for any $\mathrm{A}$ in $\mathcal{\mathrm{A}}_{n}$,
$\sup_{\mathrm{A}\in\mathcal{\mathrm{A}}_{n}}\|\mathrm{A}\phi\|^{2}\leq C_{A}\|\phi\|^{2}.$ (E.1)
Assumption E.3 (Regularity conditions for the search set $\mathcal{V}_{n}$).
I) $\left\|f\right\|_{\infty}\leq 1$ for all $f\in{\mathrm{star}}(\mathcal{V}_{n}-V_{0})$ and II) There exists a sequence $\delta_{n}=o(1)$ that is an upper bound on the critical radius of $\mathrm{star}(\mathcal{V}_{n}-V_{0})$.
Assumption E.3 is a standard regularity condition that depends on the class of estimators of the value function (see, e.g., Assumption 2 in Chernozhukov et al. (2024)).
Theorem E.1.
If Assumptions E.1 and E.3 and E.2 are satisfied, then, with probability $1-\iota$, for some constant $C_{V}$ large enough
$\left\|\widehat{V}-V_{0}\right\|^{2}\leq C_{V}\left(\delta^{2}_{n}+a^{2}_{n}+\zeta^{2}_{n}+\left\|V^{\ast}-V_{0}\right\|^{2}+\frac{\ln(1/\iota)}{n}\right),$
where $V^{\ast}$ is the best approximation of $V_{0}$ by an element of $\mathcal{V}_{n}$, that is,
$V^{\ast}=\arg\min_{V\in\mathcal{V}_{n}}\mathrm{E}[\ell_{V}(Z,V,\omega_{0})].$
Theorem E.1 gives a convergence rate in terms of the critical radius and the first-stage rate of plug-in estimators. To make use of this Theorem, one needs to know the size of the critical radius $\delta_{n}$ and the rate of approximation error $\left\|V^{\ast}-V_{0}\right\|^{2}$. For the case of MLP neural network, those have been verified in Chernozhukov et al. (2024).
Corollary E.1.
If (i) the support of $X$ is contained in a Cartesian product of compact intervals and $V_{0}(X)$ can be extended to a function that is continuously differentiable with $\beta_{V}$ continuous derivatives; (ii) $\mathcal{V}_{n}$ is an MLP network with $d_{V}$ inputs, width $K_{V}$, and depth $m_{V}$ with $K_{V}\to\infty$ and $m_{V}\to\infty$; (iii) for any estimate $\mathrm{A}\in\mathcal{A}_{n}$, $(\mathrm{A}V)_{V\in\mathcal{V}_{n}}$ is representable as such a network; then there is $C>0$ such that, for any $\varepsilon>0$,
$\left\|\widehat{V}-V_{0}\right\|^{2}=O_{p}(K_{V}^{2}m_{V}^{2}\ln(K_{V}^{2}m_{V})\ln(n)/n+[K_{V}m_{V}\sqrt{\ln(K^{2}_{V}m_{V})}]^{-2(\beta_{V}/d_{V})+\varepsilon_{V}}+\zeta^{2}_{n}+a^{2}_{n}).$
Define
$\mathrm{star}(\mathcal{A}_{n}-\alpha_{0})=~{}$$\{x\rightarrow t\,(\alpha(x)-\alpha_{0}(x)):\alpha\in\mathcal{A}_{n},\text{ }t\in[0,1]\}.$
and
$\mathrm{star}(m\circ\mathcal{A}_{n}-m\circ\alpha_{0})=~{}$$\{x\rightarrow t\,(m(z,\alpha)-m(z,\alpha_{0})):\alpha\in\mathcal{A}_{n},\text{ }t\in[0,1]\}.$
Assumption E.4.
(A) For some sequence $a^{*}_{n}=o(1)$ we have $\|\widehat{A}^{*}-A^{*}_{0}\|=O_{P}(a^{*}_{n})$.
(B) There exists a sequence $\delta_{n}=o(1)$ that is an upper bound on the critical radius of $\mathrm{star}(\mathcal{A}_{n}-\alpha_{0})$ and $\mathrm{star}(m\circ\mathcal{A}_{n}-m\circ\alpha_{0})$. (C) The estimated operator $\widehat{A}^{*}$ belongs to a realization set with probability $1-o(1)$. Each element of this realization set is mean square continuous, that is, (E.1) holds for $\mathrm{A}^{*}$ for some absolute constant $C^{*}_{A}$.
Theorem E.2.
If Assumptions C.5(C) with $\mathrm{A}^{*}$ in place of $\mathrm{A}$ and E.4 are satisfied, then, with probability $1-\iota$, for some $C_{\alpha}$ large enough
$\left\|\widehat{\alpha}-\alpha_{0}\right\|^{2}\leq C_{\alpha}\left(\delta^{2}_{n}+(a^{*})^{2}_{n}+\left\|\alpha^{\ast}-\alpha_{0}\right\|^{2}+\frac{\ln(1/\iota)}{n}\right).$
where $\alpha^{\ast}$ is the best approximation of $\alpha_{0}$ by an element of $\mathcal{A}_{n}$
$\alpha^{\ast}=\arg\min_{\alpha\in\mathcal{A}_{n}}\mathrm{E}[\ell_{\alpha}(Z,\alpha,\mathrm{A}^{*}_{0})].$
E.1. Proofs of Results of Section E
Proof of Theorem E.1.
The statement Theorem E.1 is a special case of Theorem D.1 with
$\omega=(\zeta,\mathrm{A}),\quad\ell(z,V,\omega)=\ell_{V}(Z,V,\omega).$
Steps 1, 2,3 verify Assumption D.2 (1)–(3), respectively.
Step 1. Consider $\bar{\lambda}=(1+\beta)^{2}\geq\underline{\lambda}=(1-\beta)^{2}>0$. Proposition 7.1 gives
$L(V,\omega_{0})=\|(\mathrm{I}-\mathrm{A}_{0})V-\zeta_{0}\|^{2}.$
Taking $\Delta_{V}(X)=V(X)-V_{0}(X)$ gives
$L(V,\omega_{0})-L(V_{0},\omega_{0})$ $=\|(\mathrm{I}-\mathrm{A}_{0})\Delta_{V}\|^{2}.$ (E.2)
Stationarity implies $\|\mathrm{A}_{0}\Delta_{V}\|\leq\beta\|\Delta_{V}\|$. Thus, the lower bound (D.3) follows from
$\|\Delta_{V}\|(1-\beta)=\|\Delta_{V}\|-\beta\|\Delta_{V}\|$ $\leq\|\Delta_{V}\|-\|\mathrm{A}_{0}\Delta_{V}\|\leq\|(\mathrm{I}-\mathrm{A}_{0})\Delta_{V}\|,$
and the upper bound follows from
$\|(\mathrm{I}-\mathrm{A}_{0})\Delta_{V}\|\leq\|\mathrm{I}-\mathrm{A}_{0}\|\|\Delta_{V}\|\leq(\|\mathrm{I}\|+\|\mathrm{A}_{0}\|)\|\Delta_{V}\|\leq(1+\beta)\|\Delta_{V}\|.$
Step 2. We verify Assumption D.2(2) with $C_{L}=3(1+\beta)C+2(1+\beta)$ where here $C$ is an upper bound on $\|\Delta_{V}\|\leq\|V_{0}\|\leq C$. Let $\Delta_{\zeta}=\zeta-\zeta_{0}$ and $\Delta_{\mathrm{A}}=\mathrm{A}-\mathrm{A}_{0}$. For any $V\in\mathcal{V}_{n}$, note that
$L(V,\omega)-L(V,\omega_{0})$ $=\mathrm{E}[((\mathrm{I}-\mathrm{A}_{0})V)((\Delta_{\mathrm{A}})V-2(\Delta_{\zeta}))]$
$L(V_{0},\omega)-L(V_{0},\omega_{0})$ $=\mathrm{E}[((\mathrm{I}-\mathrm{A}_{0})V_{0})((\Delta_{\mathrm{A}})V_{0}-2(\Delta_{\zeta}))].$
Define the error terms
$K_{1}$ $=\mathrm{E}[((\mathrm{I}-\mathrm{A}_{0})\Delta_{V})\cdot((\Delta_{\mathrm{A}})\Delta_{V}]$
$K_{2}$ $=\mathrm{E}[((\mathrm{I}-\mathrm{A}_{0})\Delta_{V})\cdot((\Delta_{\mathrm{A}})V_{0}]$
$K_{3}$ $=\mathrm{E}[((\mathrm{I}-\mathrm{A}_{0})V_{0})\cdot((\Delta_{\mathrm{A}})\Delta_{V}]$
$K_{4}$ $=-2\mathrm{E}[(\mathrm{I}-\mathrm{A}_{0})\Delta_{V}\cdot(\Delta_{\zeta}))].$
and note that
$L(V,\omega)-L(V,\omega_{0})-(L(V_{0},\omega)-L(V_{0},\omega_{0}))$ $=\sum_{j=1}^{4}K_{j}$
Cauchy inequality implies an upper bound
$|K_{4}|\leq 2\|(\mathrm{I}-\mathrm{A}_{0})\Delta_{V}\|\|\Delta_{\zeta}\|\leq 2(1+\beta)\|\Delta_{\zeta}\|\|\Delta_{V}\|,$
and
$\max_{j\in\{1,2,3\}}|K_{j}|\leq(1+\beta)C\|\Delta_{\mathrm{A}}\|\|\Delta_{V}\|.$
Step 3. We verify Assumption D.2(3) with $\ell_{V}(z,V,\omega)=f_{V}(V(X),V(X_{+}),\mathrm{A}V)$ where
$f_{V}(t_{1},t_{2},t_{3})=(t_{1}-\beta t_{2})(t_{1}t_{3}-\beta t_{2}t_{3})-2(t_{1}-\beta t_{2})\zeta$
that is Lipshitz with respect to the vector $(V(X),V(X_{+}),\mathrm{A}V(X))$. For any $\mathrm{A}\in\mathcal{A}_{n}$ and $\zeta$, Lemma 11 from Foster and Syrgkanis (2019) gives
$|(L_{n}(V,\omega)-L_{n}(V^{\ast},\omega))-(L(V,\omega)-L(V^{\ast},\omega))|$
$\leq^{i}O(\delta(\sqrt{\|V(X)-V_{0}(X)\|^{2}}+\sqrt{\beta^{2}\|V(X_{+})-V_{0}(X_{+})\|^{2}}+\sqrt{C\|\mathrm{A}V-\mathrm{A}V_{0}\|^{2}})+\delta^{2})$
$\leq^{ii}O(\delta\|V-V_{0}\|(1+\beta)+\sqrt{C\|\mathrm{A}V-\mathrm{A}V_{0}\|^{2}})+\delta^{2})$
$\leq^{iii}O(\delta\|V-V_{0}\|(1+\beta+\sqrt{CC_{A}})+\delta^{2})$
where (i) follows from Lemma 11, (ii) follows from stationarity and $\|\Delta_{V}(X_{+})\|=\|\Delta_{V}(X)\|$ and (iii) from MSE continuity of $\mathrm{A}$ stated in Assumption E.2.
$\blacksquare$
Proof of Theorem E.2.
The statement Theorem E.1 is a special case of Theorem D.1 with
$\omega=\mathrm{A}^{*},\quad\ell(z,V,\mathrm{A}^{*})=\ell_{\alpha}(Z,V,\mathrm{A}^{*}).$
Thus, it suffices to verify Assumption D.2 (1)–(3).
Step 1. Stationarity implies $\|\mathrm{A}_{0}\Delta_{\alpha}\|\leq\beta\|\Delta_{\alpha}\|$. An argument similar to Step 1 in the proof of Theorem E.1 verifies Assumption D.2(1) with $\bar{\lambda}=(1+\beta)^{2}\geq\underline{\lambda}=(1-\beta)^{2}>0$.
Step 2. We verify Assumption D.2(2). For any $\alpha\in\mathcal{A}_{n}$, note that
$L(\alpha,\mathrm{A}^{*})-L(\alpha,\mathrm{A}^{*}_{0})$ $=\mathrm{E}[((\mathrm{I}-\mathrm{A}^{*}_{0})\alpha)(\mathrm{A}^{*}-\mathrm{A}^{*}_{0})\alpha-2m(Z,(\mathrm{A}^{*}-\mathrm{A}^{*}_{0})\alpha)]$
$L(\alpha_{0},\mathrm{A}^{*})-L(\alpha_{0},\mathrm{A}^{*}_{0})$ $=\mathrm{E}[((\mathrm{I}-\mathrm{A}^{*}_{0})\alpha_{0})(\mathrm{A}^{*}-\mathrm{A}^{*}_{0})\alpha_{0}-2m(Z,(\mathrm{A}^{*}-\mathrm{A}^{*}_{0})\alpha_{0})]$
Subtracting the second line from the first one gives
$L(\alpha,\mathrm{A}^{*})-L(\alpha,\mathrm{A}^{*}_{0})-(L(\alpha_{0},\mathrm{A}^{*})-L(\alpha_{0},\mathrm{A}^{*}_{0}))$ $=\sum_{j=1}^{4}K^{*}_{j}$
where
$K^{*}_{1}$ $=\mathrm{E}[((\mathrm{I}-\mathrm{A}^{*}_{0})(\alpha-\alpha_{0}))\cdot(\mathrm{A}^{*}-\mathrm{A}^{*}_{0})(\alpha-\alpha_{0})]$
$K^{*}_{2}$ $=\mathrm{E}[((\mathrm{I}-\mathrm{A}^{*}_{0})(\alpha-\alpha_{0}))\cdot(\mathrm{A}^{*}-\mathrm{A}^{*}_{0})\alpha_{0}]$
$K^{*}_{3}$ $=\mathrm{E}[((\mathrm{I}-\mathrm{A}^{*}_{0})\alpha_{0})\cdot(\mathrm{A}^{*}-\mathrm{A}^{*}_{0})(\alpha-\alpha_{0})]$
$K^{*}_{4}$ $=-2\mathrm{E}[m(Z,(\mathrm{A}^{*}-\mathrm{A}^{*}_{0})(\alpha-\alpha_{0}))].$
Invoking Riesz reprepresentation (2.11) gives
$K^{*}_{4}=-2\mathrm{E}[w_{0}(X)(\mathrm{A}^{*}-\mathrm{A}^{*}_{0})(\alpha-\alpha_{0})(X)].$
Invoking Cauchy Schwartz gives an upper bound
$|K^{*}_{4}|\leq 2\|w_{0}\|\|\mathrm{A}^{*}-\mathrm{A}^{*}_{0}\|\|\alpha-\alpha_{0}\|.$
For the remaining terms, assuming $\|\alpha-\alpha_{0}\|\leq\|\alpha_{0}\|\leq C$ gives
$\max_{j\in\{1,2,3\}}|K^{*}_{j}|$ $\leq\|(\mathrm{I}-\mathrm{A}^{*}_{0})\|\|\alpha_{0}\|\|\mathrm{A}^{*}-\mathrm{A}^{*}_{0}\|\|\alpha-\alpha_{0}\|$
$\leq(1+\beta)\|\alpha_{0}\|\|\mathrm{A}^{*}-\mathrm{A}^{*}_{0}\|\|\alpha-\alpha_{0}\|.$
Step 3. We verify Assumption D.2(3) with
$f_{\alpha}(t_{1},t_{2},t_{3},t_{4})=(t_{1}-\beta t_{2})(t_{1}t_{3}-\beta t_{2}t_{3})-2t_{4}$
that is Lipshitz with respect to the vector $(\alpha(X),\alpha(X_{-}),\mathrm{A}^{*}\alpha,m(Z,(\mathrm{I}-\mathrm{A}^{*})\alpha))$. For any $\mathrm{A}^{*}\in\mathcal{A}_{n}$, Lemma 11 from Foster and Syrgkanis (2019) gives
$|(L_{n}(\alpha,\mathrm{A}^{*})-L_{n}(\alpha^{\ast},\mathrm{A}^{*}))-(L(V,\mathrm{A})-L(V^{\ast},\mathrm{A}^{*}))|$
$\leq^{i}O((\delta(\|\alpha-\alpha_{0}\|(1+\beta)+\sqrt{\|m(Z,(\mathrm{I}-\mathrm{A}^{*})\Delta_{\alpha})\|^{2}})+\delta^{2})$
$\leq^{ii}O((\delta(\|\alpha-\alpha_{0}\|(1+\beta))+\delta^{2})$
where (i) follows from invoking Lemma 11 in Foster and Syrgkanis (2019) and stationarity and (ii) from Assumption C.5(C) as well as from MSE continuity of $\mathrm{A}^{*}$ assumed in the analog of Assumption E.2.
$\blacksquare$
Appendix F Extension to Nonlinear Models
In this Section, we generalize the results of Sections 26 to allow the value function $V$ to appear nonlinearly in the model. Consider a semiparametric moment problem
$\mathrm{E}[m(Z,\delta_{0},V_{0})]=0,$ (F.1)
where $Z\in\mathcal{Z}\subseteq\mathrm{R}^{Z}$ is the data vector, $\delta$ is a finite-dimensional target parameter whose true value is $\delta_{0}$, and $V_{0}$ is value function defined in (2.1). A special case of this problem is considered in (2.9) where $m(Z,V)$ is assumed linear in $V$. In this Section, we allow both $\delta_{0}$ and $V_{0}$ to appear nonlinearly in the model.
The problem above can be translated into the theoretical framework of Section 2 via linearization of the moment function with respect to $V$, following Newey (1994). Let $D(Z,V)$ denote a function of $Z$ and a linear function $V(\cdot)$ (i.e. $D(Z,V)$ is a linear functional of $V$.) We assume that for all $V$ such that $\|V-V_{0}\|$ small enough, for some positive finite constant $\bar{B}<\infty$,
$\|m(Z,\delta_{0},V)-m(Z,\delta_{0},V_{0})-D(Z,V-V_{0})\|\leq\bar{B}\|V-V_{0}\|^{2}.$ (F.2)
We will impose throughout that the expectation $\mathrm{E}[D(Z,V)]$ is mean square continuous as a function of $V$, meaning that there is a constant $C$ such that for all $V(X)$ with $\mathrm{E}[V(X)^{2}]<\infty$,
$|\mathrm{E}[D(Z,V)]|\leq C(\mathrm{E}[V(X)^{2}])^{1/2},$ (F.3)
which is equivalent to existence of a function $w_{D0}(X)$ with $\mathrm{E}[w_{D0}(X)^{2}]<\infty$ such that
$\mathrm{E}[D(Z,V)]=\mathrm{E}[w_{D0}(X)V(X)],$ (F.4)
for all $V(\cdot)$ with $\mathrm{E}[V(X)^{2}]<\infty$ (see also e.g., Newey (1994)). In contrast to (2.11), the LHS of (F.4) involves the derivative functional rather than the original functional itself.
Let $\alpha_{0}$ be the dynamic Riesz representer corresponding to $w_{D0}$ in (F.4). The following extremum characterization applies for the $M$-estimator loss function
$\ell^{D}(Z,\alpha,\omega)=(\alpha(X)-\beta\alpha(X_{-}))((\mathrm{I}-\mathrm{A}^{*})\alpha)(X)-2D(Z,(\mathrm{I}-\mathrm{A}^{*})\alpha),\quad\omega=(\mathrm{A}^{*},D).$ (F.5)
where true value of $\omega$ is $\omega_{0}=(\mathrm{A}^{*}_{0},D_{0})$. When $\beta=0$, the Proposition F.1 recovers the extremum representation of Chernozhukov et al. (2024).
Proposition F.1 (Extremum Representation of $\alpha_{0}$ in Nonlinear Case).
The dynamic Riesz representer $\alpha_{0}$ in (5.2) is the unique minimizer of the quadratic objective function
$\alpha_{0}$ $=\arg\min_{\alpha}\mathrm{E}[(\alpha(X)-\beta\alpha(X_{-}))((\mathrm{I}-\mathrm{A}^{*})\alpha)(X)-2D(Z,(\mathrm{I}-\mathrm{A}^{*})\alpha)].$ (F.6)
$=\arg\min_{\alpha}\mathrm{E}[((\mathrm{I}-\mathrm{A}^{*})\alpha)(X)^{2}-2((\mathrm{I}-\mathrm{A}^{*})\alpha)(X)w_{D0}(X)]$
$=\arg\min_{\alpha}\mathrm{E}[(((\mathrm{I}-\mathrm{A}^{*})\alpha)(X)-w_{D0}(X))^{2}].$
Finally, let us describe the construction of the orthogonal moment for $\delta_{0}$. Let $D(Z,V,F)$ denote the plim of the estimated $D(Z,V)$ function, $\phi_{m}(Z)$ the influence function of $\mathrm{E}[D(Z,V_{0},F)]$, and $\phi_{\zeta}(Z)$ the influence function of $\mathrm{E}[\alpha_{0}(X)\zeta(X,F)]$. The orthogonal moment is
$\psi(Z,\gamma,\phi,\delta)$ $=m(Z,\delta,V)+\alpha(X)(\beta V(X_{+})-V(X)+\zeta(X))+\phi_{m}(Z)+\phi_{\zeta}(Z),$ (F.7)
$\gamma$ $=(m,V,\zeta),\phi=(\alpha,\phi_{m},\phi_{\zeta}).$
Remark F.1 (Riesz representation of Conditional Expectations of $V(X)$).
For any mean square integrable function $\rho(z)\in\mathcal{L}_{2}$, the linear functional
$D_{\rho}(Z,V):=\mathrm{E}[\rho(Z)V(X_{+})\mid X]$ (F.8)
is also mean square integrable. Its Riesz representation takes the form
$\mathrm{E}[D_{\rho}(Z,V)]=\mathrm{E}[w_{D0}(X_{+})V(X_{+})]=\mathrm{E}[w_{D0}(X)V(X)],\quad w_{D0}(X_{+}):=\mathrm{E}[\rho(Z)\mid X_{+}].$ (F.9)
Appendix G Dynamic binary choice revisited
In this Section, we demonstrate the application of Appendix F for a single-agent dynamic binary choice model obeying Assumptions 1 and 2 in Aguirregabiria and Mira (2002), described herein. At each period $t\in\{0,1,\dots,\infty\}$, the agent chooses an action $j$ in a binary set $\mathcal{A}=\{1,2\}$ after observing the state realization $x$ and the vector of private shocks $\epsilon=(\epsilon(j))_{j\in\mathcal{A}}$ independent of states. The shocks $\epsilon(j)$ are independent across $\mathcal{A}$ and type 1 extreme value distributed. For each action $j$, the state process $(X_{t}\mid X_{t-1},A_{t}=j)_{t\geq 1}$ is a first-order time-homogeneous Markov chain. The action $j$ results in $u(x,j)+\epsilon(j)$ immediate (i.e., per-period) utility. The agent chooses actions according to the optimal Markov policy
$\arg\max_{j\in\mathcal{A}}\{v(x,j)+\epsilon(j)\},$ (G.1)
where $v(x,j)$ is the choice-specific value function defined as
$v(x,1)$ $=u(x,1)+\beta\mathrm{E}[V_{0}(X_{+})\mid X=x,J=1]$ (G.2)
$v(x,0)$ $=u(x,0)+\beta\mathrm{E}[V_{0}(X_{+})\mid X=x,J=0],$ (G.3)
and $V_{0}(x)$ is the agent’s value function, determined as a fixed point of the Bellman equation
$V_{0}(x)$ $=\mathrm{E}_{\epsilon}\big{[}\max_{j\in\mathcal{A}}\big{[}v(x,j)+\epsilon(j)\mid X=x\big{]}\big{]}.$ (G.4)
As demonstrated in Hotz and Miller (1993) and Aguirregabiria and Mira (2002), $V_{0}(x)$ can be represented as a net present value (2.1) based on the per-period utility $\zeta_{0}(x)$ that we introduced in (2.4), Section 2. Furthermore, the log-odds ratio
$\ln p_{0}(x)-\ln(1-p_{0}(x))$ $=v(x,1)-v(x,0)$ (G.5)
simplifies to
$v(x,1)-v(x,0)=D(x)^{\prime}\theta_{0}+\beta(\mathrm{E}[V_{0}(X_{+})\mid X=x,J=1]-\mathrm{E}[V_{0}(X_{+})\mid X=x,J=0])$ (G.6)
for linear index utilities as in (A.3). Introducing expectation functions for the choices $J=1$ and $J=0$
$V\rightarrow\gamma^{2}_{0}(x,V)$ $=\mathrm{E}[V(X_{+})\mid X=x,J=1],$ (G.7)
$V\rightarrow\gamma^{1}_{0}(x,V)$ $=\mathrm{E}[V(X_{+})\mid X=x,J=0].$ (G.8)
gives a succinct form of log-odds ratio equality
$v(x,1)-v(x,0)=D(x)^{\prime}\theta_{0}+\beta(\gamma^{2}_{0}(x,V_{0})-\gamma^{2}_{1}(x,V_{0})).$ (G.9)
Applying the logistic transformation to (G.6) gives
$\mathrm{E}[J-\Lambda(D(x)^{\prime}\theta_{0}+\beta(\gamma^{2}_{0}(x,V_{0})-\gamma^{2}_{1}(x,V_{0}))\mid X=x]=0,$ (G.10)
where $\Lambda(t)$ and $\Lambda^{\prime}(t)$ denote the logistic CDF and the logistic PDF, specifically
$\Lambda(t)=\dfrac{\exp t}{\exp t+1},\qquad\Lambda^{\prime}(t)=\Lambda(t)\cdot(1-\Lambda(t)).$ (G.11)
For this approach to yield constructive identification, a separate identification argument for value function $V_{0}(x)$ must be provided. For instance, if the choice $J=0$ has a renewal choice property such a replacement action in bus replacement problem Rust (1987) or sterilization choice Hotz and Miller (1993), $V_{0}(x)=C-\ln(1-p_{0}(x))$ for some $C$ is identified (Example 3, Chernozhukov et al. (2022)).
G.1. Identification of $\theta_{0}$.
In what follows, let us provide a missing argument for identification of structural parameter does not require $J=0$ to be replacement, a question studied in Adusumilli and Eckardt (2019). The first part of the argument exploits linearity of per-period utility $u(x,1)$ and $u(x,0)$ in $\theta$ and is standard in the literature. It shows that value function $V_{0}(x)$ can be expressed as a linear function in $\theta$. Define the “slope” function as
$\zeta^{S}_{0}(x):=(-(1-p_{0}(x)),D_{1}(x)p_{0}(x))^{\prime}$
and the “intercept” function as
$\zeta^{I}_{0}(x):=H(p_{0}(x)),\quad H(t)=\gamma_{e}-t\ln t-(1-t)\ln(1-t).$
Next, consider a linear form of $\theta$
$\zeta(x,\theta)=\zeta^{S}_{0}(x)^{\prime}\theta+\zeta^{I}_{0}(x),$ (G.12)
which coincides with the per-period utility $\zeta_{0}(x)$ in (2.4) at the true value $\theta=\theta_{0}$. Likewise, the linear form
$V(x,\theta)=:V^{S}_{0}(x)^{\prime}\theta+V^{I}_{0}(x)$ (G.13)
constructed using the slope and intercept functions
$V^{S}_{0}(x)=\sum_{t\geq 0}\beta^{t}\mathrm{E}[\zeta^{S}(X_{t})\mid X=x],\quad V^{I}_{0}(x)=\sum_{t\geq 0}\beta^{t}\mathrm{E}[\zeta^{I}(X_{t})\mid X=x]$ (G.14)
matches the value function $V_{0}(x)$ in (2.1) at at the true value $\theta=\theta_{0}$. Our proposal is to replace value function $V$ in (G.10) by a linear form $V(x,\theta)$, treating $V^{S},V^{I}$ as nuisance functions. Then, (G.10) is a semiparametric conditional moment restriction whose parametric component is $\theta$ and identified functional component is $h_{0}(x)=(V^{S}_{0}(x),V^{I}_{0}(x),\gamma^{2}_{0}(x),\gamma^{1}_{0}(x))$ where the true values of expectation functions $\gamma^{1},\gamma^{2}$ are given in (G.7)–(G.8) and true values of $V^{S},V^{I}$ are given in (G.14). Taking
$\upsilon_{1}(X,h)$ $=D(x)+\beta(\gamma^{2}(x,V^{S})-\gamma^{1}(x,V^{S}))$
$\upsilon_{0}(X,h)$ $=\beta(\gamma^{2}_{0}(x,V^{I})-\gamma^{1}_{0}(x,V^{I}))$
$\upsilon(X,h)$ $=(\upsilon^{\prime}_{1}(X,h),\upsilon^{\prime}_{0}(X,h))^{\prime}$
gives
$\upsilon(x,h_{0})^{\prime}\theta_{0}=v(x,1)-v(x,0)=\ln p_{0}(x)-\ln(1-p_{0}(x)).$
Invoking logistic transform to both sides gives
$p_{0}(x)=\Lambda(\upsilon(x,h_{0})^{\prime}\theta_{0})$
which gives rise to an unconditional moment restriction
$g_{\theta}(Z,\theta,h)=\upsilon(X,h)(J-\Lambda(\upsilon(X,h)^{\prime}\theta)).$ (G.15)
G.2. Overview of estimation and inferential strategy.
Our estimation strategy for the structural parameter $\theta_{0}$ is to combine the extremum representation of $V^{S}_{0},V^{I}_{0}$ with the orthogonal moment $\psi$ which we derive below. We follow the outline of Section 6 and begin with the derivation of debiased moment function. The moment function (F.7) reduces to
$\psi(Z,h,\alpha,\theta)$ $=\upsilon(X,h)(J-\Lambda(\upsilon(X,h)^{\prime}\theta_{0}))+\sum_{j=1}^{2}\phi_{\gamma^{j}}(Z,h)$ (G.16)
$+\alpha(X)(\beta V(X_{+})-V(X)+\zeta(X))+\phi_{\zeta}(Z,h)$ (G.17)
where the first two correction terms $\phi_{\gamma^{1}}$ and $\phi_{\gamma^{2}}$ in (G.16) account for the estimation of $\gamma^{1}$ and $\gamma^{2}$, respectively and the third one in (G.17) accounts for the nuisance components of linear form $V(x,\theta)$. A standard argument for expectation functions (Newey (1994), Proposition 4, p. 1361) gives
$\phi_{\gamma^{2}}(Z,h)$ $=\alpha_{2}(X,J)(V(X_{+})-\gamma_{2}(X,V))$ (G.18)
$\phi_{\gamma^{1}}(Z,h)$ $=\alpha_{1}(X,J)(V(X_{+})-\gamma_{1}(X,V)),$ (G.19)
where the true values of bias correction functions $\alpha_{2}(X,J)$ and $\alpha_{1}(X,J)$ are
$\alpha_{20}(X,J)$ $=-\beta\upsilon(X)J(1-p_{0}(X))$ (G.20)
$\alpha_{10}(X,J)$ $=\beta\upsilon(X)(1-J)p_{0}(X).$
The third correction term in line (G.17) is specified up to $\alpha$, which can be derived as described in Section F, and the correction term $\phi_{\zeta}$. Note that the linearization functional (F.2)
$D(Z,V):$ $=-\beta\cdot\upsilon(X)\cdot p_{0}(X)\cdot(1-p_{0}(X))\cdot\left(\gamma^{2}_{0}(X,V)-\gamma^{1}_{0}(X,V)\right)$ (G.21)
$=-\beta\mathrm{E}[(\alpha_{2}(X,J)-\alpha_{1}(X,J))V(X_{+})\mid X]$
is a special case of a linear functional (F.8) in Remark F.1 with $\rho(Z)=\alpha_{2}(X,J)-\alpha_{1}(X,J)$. Thus, the true value of $\alpha$ is the dynamic Riesz representer based on weighting function
$w_{D0}(x)$ $=-\beta\mathrm{E}[\upsilon(X)(J-p(X))\mid X_{+}=x]$ (G.22)
whose automatic characterization is given in Proposition F.1 with $D(Z,V)$ in (G.21). Finally, the fourth correction term accounts for estimation of choice probability $p(x)$ only  [^6] This approach is different from Section A where both $p_{0}(x)$ and $\theta$ are treated as nuisance components. that enters linearly in $\zeta^{S}(x)$ and nonlinearly in $\zeta^{I}(x)=H(p(x))$. The correction term is
$\phi_{\zeta}(Z,h)$ $=-\beta\alpha(X)(\gamma^{2}(X,V)-\gamma^{1}(X,V))(J-p(X)).$ (G.23)
Remark G.1 (Static models with $\beta=0$).
Notice all correction terms $\phi_{\gamma^{2}}(Z,h),\phi_{\gamma^{1}}(Z,h),\phi_{\zeta}(Z,h)$ and $\alpha(x)$ in (G.20)–(G.23) depend multiplicatively on $\beta$ and, therefore, reduce to zero if $\beta=0$. In this case, the moment equation (G.15) reduces to a FOC for the logistic regression
$\mathrm{E}[D(X)(J-\Lambda(D(X)^{\prime}\theta_{0}))]=0,$
which has no nuisance parameters. The same property hold for the debiased moment condition in Example 3, Chernozhukov et al. (2022).
Recent work by Adusumilli and Eckardt (2019) considered semiparametric estimation and inference method for a structural parameter in dynamic discrete choice models. The approach proposed in Appendix G complements their results in several respects. First, the authors consider data arriving in state-action pairs $((X,J),(X_{+},J_{+}))$ while this paper considers standard triplets $(X,J,X_{+})$, following Aguirregabiria and Mira (2002) and Chernozhukov et al. (2022). Consequently, the nuisance parameters of the moment condition proposed therein are based on choice-specific value functions $(v(x,j))_{j\in\mathcal{A}}$, while the nuisance functions of (G.14)–(G.15) are slope and intercept of value function $V_{0}(x)$.
Algorithm 2 sketches the proposed estimator of $\theta$.
1:  Partition the set of data indices $\{1,2,\dots,n\}$ into $L$ disjoint subsets of about equal size with where $L$ is an odd number $L\geq 3$. 2:  Let $I^{c}_{\ell}=(Z_{i})_{i\notin I_{\ell}}$ denote the set observations not in $I_{\ell}$. Partition $I^{c}_{\ell}=I^{1}_{\ell}\sqcup I^{2}_{\ell}$ into two halves. For each nuisance parameter $\omega$, let $\widehat{\omega}^{1}_{\ell},\omega^{2}_{\ell}$ denote the estimator computed on $I^{1}_{\ell}$ and $I^{2}_{\ell}$, respectively. Estimate a The choice probability function $p_{0}(x)={\mathrm{P}}(J=1\mid X=x)$ b The expectation functions $V\rightarrow\gamma^{1}(x,V)$ and $V\rightarrow\gamma^{2}(x,V)$ as in (G.8) and (G.7). c The forward operator $\mathrm{A}$ in (7.1) and backward operator $\mathrm{A}^{*}$ in (7.15) 3:  Estimate “slope” and “intercept” value functions by minimizing sample cross-fit objective function $\widehat{V}^{S}_{\ell}=\arg\min_{V\in\mathcal{V}_{n}}\bigg{[}\sum_{i\in I^{c1}_{\ell}}\ell^{S}(Z_{i},\widehat{\omega}^{2}_{\ell})+\sum_{i\in I^{c2}_{\ell}}\ell^{S}(Z_{i},\widehat{\omega}^{1}_{\ell})\bigg{]},\qquad\widehat{\omega}=(\widehat{\mathrm{A}},\widehat{p})$ $\widehat{V}^{I}_{\ell}=\arg\min_{V\in\mathcal{V}_{n}}\bigg{[}\sum_{i\in I^{c1}_{\ell}}\ell^{I}(Z_{i},\widehat{\omega}^{2}_{\ell})+\sum_{i\in I^{c2}_{\ell}}\ell^{I}(Z_{i},\widehat{\omega}^{1}_{\ell})\bigg{]},\qquad\widehat{\omega}=(\widehat{\mathrm{A}},\widehat{p})$ where $\mathcal{V}_{n}$ is some set of functions and $\ell^{S}$ and $\ell^{I}$ are special cases of $M$-estimator losses in (7.8) based on $\zeta^{S}(x)$ and $\zeta^{I}(x)$, respectively. 4:  Estimate dynamic Riesz representer by minimizing sample cross-fit objective function $\widehat{\alpha}_{\ell}=\arg\min_{\alpha\in\mathcal{A}_{n}}\bigg{[}\sum_{i\in I^{c1}_{\ell}}\ell^{D}(Z_{i},\widehat{p}_{\ell},\widehat{\gamma}^{2}_{\ell},\widehat{\gamma}^{1}_{\ell},\widehat{\mathrm{A}}^{2*}_{\ell})+\sum_{i\in I^{c2}_{\ell}}\ell^{D}(Z_{i},\widehat{p}_{\ell},\widehat{\gamma}^{2}_{\ell},\widehat{\gamma}^{1}_{\ell},\widehat{\mathrm{A}}^{1*}_{\ell})\bigg{]},$ where $\mathcal{A}_{n}$ is some set of functions and $\ell^{D}$ is a special case of $M$-estimator losses in (F.5) based on $D(Z,V)$. 5:  Estimate the nuisance parameters $(\widehat{p}_{\ell},\widehat{\gamma}^{2}_{\ell},\widehat{\gamma}^{1}_{\ell})$ using all observations in $I^{c}_{\ell}$. Define a collection of estimated functions as $\widehat{h}_{\ell}=(\widehat{p}_{\ell},\widehat{\gamma}^{2}_{\ell},\widehat{\gamma}^{1}_{\ell},\widehat{(}V^{S}_{\ell})^{\prime}\widehat{\theta}_{\ell}+\widehat{V}^{I}_{\ell},\widehat{(}\zeta^{S}_{\ell})^{\prime}\widehat{\theta}_{\ell}+\widehat{\zeta}^{I}_{\ell})$ where $\widehat{\theta}_{\ell}$ is a preliminary estimator of $\theta$ based on a non-orthogonal moment condition. 6:  Estimate the structural parameter using debiased GMM estimator $\frac{1}{n}\sum_{\ell=1}^{L}\sum_{i\in I_{\ell}}\psi(Z_{i},\widehat{\theta},\widehat{h}_{\ell})=0$
Algorithm 2 Debiased Estimator of Structural Parameter
G.3. Auxiliary Empirical Details
Table 4. Descriptive Statistics of Key Variables
Full Low Medium High
Sample Score $\in[3,30]$ Score $\in[30,40]$ Score $\in[40,54]$
School Open 0.78 0.73 0.80 0.80
(0.41) (0.45) (0.40) (0.40)
Teacher Test Scores 35.08 20.22 36.61 43.92
(10.68) (6.89) (2.33) (3.67)
Days Worked 17.15 16.07 17.64 17.50
(5.55) (6.29) (5.18) (5.18)
Work Streak 3.61 3.10 3.73 3.86
(5.55) (4.94) (5.68) (5.82)
Number of Observations 19016 5245 5984 7787
Number of Teachers 57 16 17 24
Notes: This table reports sample means and standard deviations of School Open, Days Worked, Work Streak and Teacher Test Score where the latter variable is used to categorize teachers into Low-, Medium-, and High-score groups. The data sample comes from Duflo et al. (2012) and spans teacher attendance data between January 1, 2004, and June 30, 2005, excluding holidays and Sundays when all schools were closed. See text for details.